### v2

# Quasi-Clique Discovery via Energy Diffusion

Abstract
Discovering quasi-cliques—subgraphs whose edge density exceeds a given threshold—is a fundamental task in graph mining with applications to web spam detection, fraud screening, and e-commerce recommendation. However, existing methods for quasi-clique discovery on large-scale web graphs are often sensitive to random seeds or lack of explicit edge-density guarantees, making the task challenging in practice. This paper presents EDQC, an energy diffusion–based method for quasi-clique discovery. EDQC first employs an adaptive energy diffusion process to generate an energy ranking that highlights structurally cohesive regions. Guided by this energy ranking, the algorithm identifies a high-quality subgraph by minimizing conductance, a standard measure from community detection. This subgraph is then refined to meet the specified density threshold. Extensive experiments on 75 real-world graphs show that EDQC finds larger quasi-cliques on most datasets, with consistently lower variance across runs and competitive runtime. To the best of our knowledge, EDQC is the first method to incorporate energy diffusion into quasi-clique discovery.
1 Introduction
Dense subgraph models are fundamental to graph mining on large-scale web graphs, enabling web spam detection gibson2005discovering ; lin2021mining , fraud screening zheng2019author ; fang2022densest , and e-commerce recommendation schafer2001commerce ; song2021subgraph . Among these models, the clique represents the most stringent form, requiring that every pair of vertices be directly connected to each other saito2007large . However, due to noise and missing links in real-world graphs, exact cliques are rarely observed. To relax this strict requirement, several approximate dense subgraph models have been proposed, including quasi-cliques abello2002massive ; pei2005mining , $k$-clubs mokken1979cliques , and $k$-plexes seidman1978graph . In particular, the $\gamma$-quasi-clique, defined by a minimum edge-density threshold, is one of the most widely adopted formulations chen2021nuqclq ; pang2024similarity ; liu2024optimization .
Given an undirected graph $G$ and an edge-density threshold $\gamma\in(0,1]$, the Maximum Quasi-Clique Problem (MQCP) seeks the largest vertex set $S$ such that the induced subgraph $G[S]$ exhibits an edge density of at least $\gamma$. This edge-density criterion provides a more direct measure of structural cohesion compared to the degree-based definitions utilized by methods such as FastQC konar2024optimal and IterQC liu2008effective . However, the MQCP is NP-hard even for fixed values of $\gamma$ pattillo2013maximum and does not possess the hereditary property that allows for efficient pruning in problems related to cliques xia2025maximum . Nonetheless, quasi-clique discovery remains highly valuable, as real-world graphs typically exhibit dense yet incomplete connectivity rather than perfect cliques, making quasi-cliques crucial for identifying significant or anomalous substructures.
Figure 1: A clique and a 0.8-quasi-clique.
One practical application of quasi-clique discovery is identifying link farm spam pages in web graphs pang2024similarity , where densely connected spam pages form $\gamma$-quasi-cliques, enabling the identification of suspicious link farms even under noise and missing links. As shown in Figure 1, the dashed edges in the left subgraph form a size-4 clique, representing a fully connected link farm. In the right subgraph, the dashed edges form a 0.8-quasi-clique, still suspicious despite one missing link, capturing a potential link farm under partial observation or evasion.
Despite recent progress, discovering quasi-cliques in real large-scale web graphs remains challenging due to both scale and combinatorial complexity. While exact algorithms for MQCP, typically based on integer programming or techniques such as branch-and-bound uno2010efficient ; pastukhov2018maximum ; ribeiro2019exact ; rahman2024fast , guarantee optimal solutions, their exponential time complexity restricts their applicability to large graphs. As a practical alternative, numerous heuristic methods have been proposed to enable scalable quasi-clique discovery, which can be roughly grouped into four categories. First, greedy expansion techniques, such as the greedy strategy with local refinement introduced in tsourakakis2013denser and the restart-enabled search in RIG oliveira2013construction , prioritize speed but are vulnerable to becoming trapped in local optima. Second, metaheuristic frameworks like TSQC djeddi2019extension and OBMA zhou2020opposition employ tabu search and opposition-based learning, offering enhanced robustness at a higher computational cost. Third, local search algorithms, including NuQClq chen2021nuqclq and its variants NuQClqF1 and NuQClqR1 liu2024optimization , utilize bounded configuration checking and adaptive rules for candidate refinement. However, despite their sophisticated local optimizations, these methods often exhibit instability due to their dependence on random seeds. Lastly, similarity-based heuristics such as NBSim and its accelerated variant, FastNBSim, utilize vertex neighborhood overlap to identify and expand quasi-cliques pang2024similarity . While both methods are highly efficient, FastNBSim introduces MinHash-based approximations, making it sensitive to random seeds and potentially less stable. Moreover, neither method supports a user-specified edge-density threshold. Instead, the edge-density of the returned subgraph emerges implicitly from similarity parameters. Consequently, there is no guarantee that the output meets a specified minimum edge-density requirement. This limits their suitability for applications that require strict control over quasi-clique edge-density thresholds.
Among these four categories, local search and similarity-based methods have been the most competitive for discovering large quasi-cliques. The state-of-the-art algorithms for MQCP include NBSim and its accelerated variant, FastNBSim pang2024similarity , as well as the local search algorithms NuQClqF1 and NuQClqR1 liu2024optimization . However, these methods remain sensitive to random seeds or lack of explicit edge-density guarantees.
Energy diffusion is a natural candidate for addressing the challenges above. Diffusion methods such as PageRank page1999pagerank and the heat kernel chung2007heat have shown strong performance on structure-aware tasks, e.g., vertex ranking and community detection kloster2014heat ; gleich2015multilinear , where diffused mass tends to localize within cohesive regions. These tasks overlap with quasi-clique discovery in highlighting tightly interlinked vertices, yet they are not equivalent because quasi-cliques must satisfy a specified minimum edge-density on induced subgraphs. This connection motivates exploring energy diffusion for quasi-clique discovery to reduce sensitivity to random seeds in practice and to furnish a stable candidate region for density-constrained extraction.
This paper introduces EDQC, an energy diffusion-based method for quasi-clique discovery. EDQC begins with an adaptive energy diffusion phase, where the energy transfer rate is dynamically adjusted based on the local energy distribution. This phase runs for a fixed number of iterations from each source vertex, generating an energy ranking that highlights structurally cohesive regions. Guided by this energy ranking, EDQC then employs a principled extraction mechanism from community detection theory. It applies a sweep cut to the energy-ranked vertices to identify a high-quality subgraph that minimizes conductance. This subgraph is subsequently refined through a structure-driven procedure to guarantee it meets the specified density threshold. Extensive experiments on 75 real-world graphs, including classic web graphs, demonstrate that EDQC discovers larger quasi-cliques on most datasets, with consistently lower variance and competitive runtime. To the best of our knowledge, EDQC is the first method to incorporate energy diffusion into quasi-clique discovery.
2 Preliminaries
All graphs considered in this paper are undirected and simple (no loops or parallel edges). Let $G=(V,E)$ represent a graph with a vertex set $V$ and an edge set $E$. Two vertices $u$ and $v$ within $V$ are termed adjacent if $uv\in E$, and one vertex is referred to as a neighbor of the other. The set of all neighbors for a vertex $u$ is known as its neighborhood, denoted $N(u)$, while the closed neighborhood of $u$ is denoted $N[u]=N(u)\cup\{u\}$. The degree of a vertex $u$, denoted as $d(u)$, indicates the number of its neighbors, i.e., $d(u)=|N(u)|$.
Let $n=|V|$ and $m=|E|$. For a vertex set $S\subseteq V$, the induced subgraph is $G[S]$ with vertex set $S$ and edge set $E(S)=\{\{u,v\}\in E\mid u,v\in S\}$. Let $m(S)=|E(S)|$ denote the number of internal edges in $G[S]$, and define the internal degree of $u\in S$ by
$d_{S}(u)=|N(u)\cap S|.$
The edge density of $S$ (and of $G[S]$) is
$\eta(S)=\begin{cases}\dfrac{2\,m(S)}{|S|(|S|-1)},&|S|\geq 2,\\[6.0pt] 0,&otherwise,\end{cases}$
and $\eta(G)$ is shorthand for $\eta(V)$. Throughout this paper, unless otherwise stated, density means edge density.
Definition 1 ($\gamma$-Quasi-Clique).
Given $G=(V,E)$ and a density threshold $\gamma\in(0,1]$. A vertex set $S\subseteq V$ induces a $\gamma$-quasi-clique if the induced subgraph $G[S]$ has edge density at least $\gamma$, i.e.,
$\eta(G[S])\;\geq\;\gamma.$
For brevity, we also refer to $G[S]$ itself as a $\gamma$-quasi-clique. A $\gamma$-quasi-clique $G[S]$ is maximal if there is no $v\in V\setminus S$ such that $\eta\!\left(G[S\cup\{v\}]\right)\geq\gamma$.
Definition 2 (Maximum Quasi-Clique Problem).
Given $G=(V,E)$ and a density threshold $\gamma\in(0,1]$, the Maximum Quasi-Clique Problem (MQCP) seeks
$S^{*}\in\arg\max_{S\subseteq V}\ \Big\{|S|\ \Big|\ \eta(G[S])\geq\gamma\Big\}.$
The induced subgraph $G[S^{*}]$ is a maximum $\gamma$-quasi-clique. When $\gamma=1$, MQCP reduces to the classical Maximum Clique Problem. For any fixed $\gamma$, MQCP is NP-hard and may admit multiple optimal solutions pattillo2013maximum .
Definition 3 (Conductance).
Conductance is a standard measure used to evaluate how well a vertex set forms a community andersen2006local . For a non-empty, proper subset of vertices $S\subset V$, its volume is defined as $\mathrm{vol}(S)=\sum_{u\in S}d(u)$, and the cut between $S$ and its complement is $\mathrm{cut}(S,V\setminus S)=|\{\{u,v\}\in E\mid u\in S,v\in V\setminus S\}|$. The conductance of $S$ is then:
$\Phi(S)=\frac{\mathrm{cut}(S,V\setminus S)}{\min(\mathrm{vol}(S),\mathrm{vol}(V\setminus S))}.$
A lower conductance value indicates a more clearly defined community, signifying few edges connecting the set to the rest of the graph relative to its internal connectivity.
3 The EDQC Algorithm
This section details the proposed Energy Diffusion-based Quasi-Clique discovery (EDQC) algorithm. Given a graph $G=(V,E)$ and a density threshold $\gamma$, EDQC identifies a large vertex set $S$ that induces a $\gamma$-quasi-clique. The overall pipeline is outlined in Algorithm 1.
The algorithm iterates through vertices, treating each as a source to initiate a two-stage process. First, an adaptive energy diffusion process generates an energy distribution over the graph. Second, guided by this energy distribution, a principled extraction mechanism identifies a high-quality quasi-clique. EDQC maintains the largest quasi-clique found across all source vertices. Processing vertices in non-increasing degree order is an efficiency heuristic rossi2014fast , as it ensures high-degree vertices are processed as energy sources first, thereby increasing the likelihood of finding a large quasi-clique early in the search.
Input: Graph $G=(V,E)$, density threshold $\gamma$, diffusion rounds $T$, activation threshold $\theta$
Output: Vertex set $S^{*}$ of a large $\gamma$-quasi-clique
1 $S^{*}\leftarrow\emptyset$;
2 foreach $v\in V$ in non-increasing order of degree do
[…]
5 if $|S|>|S^{*}|$ then
6    $S^{*}\leftarrow S$;
7
8
return $S^{*}$
[…]
3.1 Adaptive Energy Diffusion
The first stage, detailed in Algorithm 2, simulates the diffusion of energy from a source vertex $v$ over $T$ synchronous rounds. The goal is to produce an energy function $f:V\to\mathbb{R}_{\geq 0}$ where energy is concentrated in structurally cohesive regions surrounding the source.
The diffusion process is designed to be adaptive. The fraction of energy a vertex releases depends on its current energy relative to its neighborhood, a dynamic governed by an adaptive fraction $\alpha(u)$. This design is deliberate and distinct from standard diffusion mechanisms such as the heat kernel chung2007heat or PageRank page1999pagerank , which often treat connections uniformly. Our adaptive fraction makes the diffusion process sensitive to the evolving energy landscape, enabling EDQC to be more conservative in sparse regions and more aggressive within dense ones.
The diffusion starts by initializing a unit energy at the source vertex $v$ (line 1). In each of the $T$ synchronous rounds (lines 2–14), the algorithm first identifies the set of active vertices, $A$, based on the activation threshold $\theta$ (line 3). To ensure updates are synchronous, a temporary copy of the energy function, $f^{\prime}$, is created (line 4). The diffusion rule is then applied to each active vertex $u\in A$ with a non-zero degree ($d(u)>0$). Isolated vertices are skipped as they have no neighbors to which energy can be transferred, a condition necessary for the conservation of total energy. For each such vertex $u\in A$ (line 5), the algorithm computes its adaptive fraction $\alpha(u)$ and the total energy to transfer, $\delta(u)$ (lines 7–8). This energy is then stochastically allocated among its neighbors using weights $\phi_{uw}$ derived from an exponential distribution (lines 9–10). Subsequently, the energies of the active vertex and its neighbors are updated in the temporary function $f^{\prime}$ (lines 11–13). After all active vertices have been processed, the energy function $f$ is updated with the temporary copy, completing the round (line 14).
Input: Graph $G=(V,E)$, source vertex $v$, diffusion rounds $T$, activation threshold $\theta$
Output: Energy function $f:V\to\mathbb{R}_{\geq 0}$
1 Initialize $f(v)\leftarrow 1$ and $f(x)\leftarrow 0$ for $x\neq v$;
2 for $t=1$ to $T$ do
3 $A\leftarrow\{x\in V\mid f(x)>\theta\}$;
4 $f^{\prime}\leftarrow f$;
5 foreach $u\in A$ do
6    if $d(u)>0$ then
7       $\alpha(u)\leftarrow\frac{\sum_{w\in N(u)}f(w)+f(u)}{\sum_{w\in N(u)}f(w)+2f(u)}$;
8       $\delta(u)\leftarrow\alpha(u)f(u)$;
9         Sample $\xi_{uw}\sim\mathrm{Exp}(1)$ for each $w\in N(u)$;
10       $\phi_{uw}\leftarrow\frac{\xi_{uw}}{\sum_{x\in N(u)}\xi_{ux}}$ for each $w\in N(u)$;
11       $f^{\prime}(u)\leftarrow f^{\prime}(u)-\delta(u)$;
12       foreach $w\in N(u)$ do
13          $f^{\prime}(w)\leftarrow f^{\prime}(w)+\delta(u)\phi_{uw}$;
14
15
16
17 $f\leftarrow f^{\prime}$;
18
return $f$
[…]
Properties.
This diffusion process has several key properties.
Proposition 1 (Energy Conservation).
The total energy is conserved in each round: $\sum_{x\in V}f^{\prime}(x)=\sum_{x\in V}f(x)$.
Proof.
Consider a fixed round and an arbitrary active vertex $u$ where $d(u)>0$. The synchronous updates associated with $u$ results in a change of $\delta(u)$ for $u$ and an addition of $\delta(u)\phi_{uw}$ for each neighbor $w\in N(u)$, satisfying $\sum_{w\in N(u)}\phi_{uw}=1$. Therefore, $\delta(u)=\sum_{w\in N(u)}\delta(u)\phi_{uw}$. Summing over all active vertices leads to the conclusion that $\sum_{x\in V}f^{\prime}(x)=\sum_{x\in V}f(x)$. Here, we should note that the energy of vertices for which $d(u)=0$ or $f(u)=0$ remains unchanged in the energy diffusion process. ∎
The conservation of energy is a crucial property for the algorithm’s stability. It guarantees that the process only redistributes the initial energy, rather than arbitrarily creating or losing it, thus ensuring a fair basis for comparing the energy distributions originating from different source vertices.
Proposition 2 (Properties of $\alpha(u)$).
For an active vertex $u$ with $f(u)>0$ and $d(u)>0$, the adaptive fraction $\alpha(u)\in[1/2,1)$ and is monotonically increasing with the total energy of its neighbors, $S(u)=\sum_{w\in N(u)}f(w)$.
Proof.
We begin by rewriting the formula for $\alpha(u)$ (line 7, Algorithm 2) as $\alpha(u)=1-\frac{f(u)}{S(u)+2f(u)}$. Since $S(u)\geq 0$ and $f(u)>0$ for an active vertex, the denominator $S(u)+2f(u)$ is strictly positive. The fractional term is bounded by $0<\frac{f(u)}{S(u)+2f(u)}\leq\frac{f(u)}{2f(u)}=1/2$. Consequently, $\alpha(u)$ is bounded by $1-1/2\leq\alpha(u)<1-0$, which gives $\alpha(u)\in[1/2,1)$. For monotonicity, the partial derivative with respect to $S(u)$ is:
$\frac{\partial\alpha(u)}{\partial S(u)}=\frac{f(u)}{(S(u)+2f(u))^{2}}>0.$
This confirms that $\alpha(u)$ is strictly increasing with $S(u)$. ∎
The adaptive nature of our diffusion is formally established by this proposition. The bounds on $\alpha(u)$ guarantee a process that is both consistently active and stable, preventing the diffusion from stalling. Concurrently, the monotonicity is the key mechanism driving the algorithm’s core premise: it mathematically ensures that energy is exchanged more readily within already high-energy (and thus likely dense) regions, accelerating the concentration of energy.
Proposition 3 (Unbiased Distribution in Expectation).
The expected fraction of energy that an active vertex $u$ transfers to a specific neighbor $w\in N(u)$ is uniform across all neighbors: $\mathbb{E}[\phi_{uw}]=1/d(u)$.
Proof.
The weights $\phi_{uw}$ are derived by normalizing i.i.d. exponential random variables $(\xi_{uw})_{w\in N(u)}$. This construction is a standard method for generating a random point from a uniform distribution on the standard simplex. By symmetry, the expectation of each component is identical. Since $\sum_{w\in N(u)}\phi_{uw}=1$, by linearity of expectation, $\sum_{w\in N(u)}\mathbb{E}[\phi_{uw}]=1$. As there are $d(u)$ identical terms in the sum, it follows that $d(u)\cdot\mathbb{E}[\phi_{uw}]=1$, which proves $\mathbb{E}[\phi_{uw}]=1/d(u)$. ∎
While the exponential sampling introduces stochasticity to explore different structural paths, this proposition guarantees that the process remains unbiased in expectation. On average, all neighbors are treated fairly, ensuring that the algorithm itself introduces no arbitrary structural bias during the diffusion.
Theorem 1 (Complexity of Algorithm 2).
The EnergyDiffusion algorithm runs in $O(T(n+m))$ time.
Proof.
The algorithm performs $T$ rounds. We analyze the computational cost of a single round. First, identifying the active set $A$ requires a full scan of all vertices, which takes $O(n)$ time. Next, the algorithm iterates through each active vertex $u\in A$. For each such vertex, all calculations—including the summations for $\alpha(u)$ and $\phi_{uw}$, and the subsequent updates—are proportional to its degree, $d(u)$. Therefore, the total work for processing all active vertices in a round is bounded by the sum of degrees of all vertices in the graph, which is $\sum_{u\in V}O(d(u))=O(m)$. Summing the costs within a single round gives a complexity of $O(n+m)$. As the process repeats for $T$ rounds, the total complexity of the algorithm is $O(T(n+m))$. ∎
3.2 Quasi-Clique Extraction
The second stage of EDQC takes the final energy function $f$ returned by Algorithm 2 and produces a vertex set $S$ that induces a $\gamma$-quasi-clique. This extraction process (Algorithm 3) has three main operations: an initial ranking of vertices, the identification of a high-quality initial subgraph, and a refinement procedure.
First, the algorithm identifies the set of active vertices $A$ and sorts them in descending order of their energy scores to create an ordered sequence $P$ (lines 1–2). This ranking is crucial, as it places vertices most likely to belong to a cohesive community at the forefront.
Next, to identify a high-quality initial candidate set, the algorithm performs a sweep cut andersen2006local on the energy-ranked list $P$ (line 3). This procedure examines the nested prefixes of the list and selects the one that minimizes conductance, as it represents the most well-defined community structure. This strategy is highly effective because the initial energy ranking provides a high-quality vertex permutation. It is a known principle in local graph clustering that sweep cuts perform best when applied to an ordering that places vertices from the same community contiguously andersen2006local . Our energy diffusion process is designed for precisely this purpose: it creates an ordering where structurally cohesive vertices are naturally placed at the top. This, in turn, enables the sweep cut to accurately identify the boundary of the best candidate subgraph.
Finally, this initial set $S$ undergoes a refinement procedure (lines 4–15). The refinement begins with a pruning loop (lines 4–6). If the initial set $S$ does not meet the density threshold, this loop iteratively removes the vertex with the minimum internal degree $d_{S}(u)$ until the set $S$ induces a valid $\gamma$-quasi-clique. Following this, an expansion phase seeks to enlarge the set (lines 8–15). In each pass, the remaining candidates $R$ are re-sorted based on their connectivity to the current set $S$ (line 10). The algorithm then iterates through this sorted list $R$, attempting to greedily add the first vertex $v$ that maintains the density constraint (lines 12–13). Upon a successful insertion, the current iteration over the candidate list is immediately terminated (line 15), allowing the algorithm to restart the process with an up-to-date candidate ranking in the next iteration.
Input: Graph $G=(V,E)$, energy function $f$, density threshold $\gamma$, activation threshold $\theta$
Output: A vertex set $S$ that induces a $\gamma$-quasi-clique
1
[…]
3 $P\leftarrow$ vertices of AA sorted by decreasing $f$;
4
[…]
6
7while $|S|>1$ and $\eta(G[S])<\gamma$ do
8 $u\leftarrow\arg\min_{x\in S}d_{S}(x)$;
9 $S\leftarrow S\setminus\{u\}$;
10
11
[…]
13 while changed do
14 $\textit{changed}\leftarrow\textbf{false}$;
15 $R\leftarrow$ vertices of $P\setminus S$, sorted by connections to $S$;
16 foreach $v\in R$ do
17    if $\eta(G[S\cup\{v\}])\geq\gamma$ then
18       $S\leftarrow S\cup\{v\}$;
19       $\textit{changed}\leftarrow\textbf{true}$;
20       break;
21
22
23
return $S$
[…]
Theorem 2 (Complexity of Algorithm 3).
Let $k=|A|$ be the number of active vertices. The ExtractQuasiClique stage runs in $O(k^{2}+m(A))$, where $m(A)$ is the number of edges in the induced subgraph $G[A]$.
Proof.
The overall complexity is determined by the main steps of the algorithm. Sorting the $k$ active vertices takes $O(k\log k)$ time. The subsequent sweep cut, which involves calculating conductance for $k$ prefixes, can be implemented in $O(m(A))$ time with incremental updates for the cut and volume. The final refinement stage is the most computationally intensive. The pruning and expansion loops can, in the worst case, lead to a series of $O(k)$ vertex additions or removals. With incremental maintenance of internal degrees and edge counts, each step of these loops can take up to $O(k)$ time to find the best candidate to remove or to scan through remaining candidates for addition. This results in a worst-case complexity of $O(k^{2})$ for the refinement phase. Summing these costs, the total complexity is $O(k\log k+m(A)+k^{2})$, which simplifies to $O(k^{2}+m(A))$. ∎
Theorem 3 (Complexity of EDQC).
The theoretical worst-case time complexity of the EDQC algorithm (Algorithm 1) is $O(n^{3})$.
Proof.
The algorithm’s main loop iterates up to $n$ times. Each iteration calls two procedures: EnergyDiffusion, with a complexity of $O(T(n+m))$, and ExtractQuasiClique, with a worst-case complexity of $O(n^{2}+m)$. As $T$ is a small constant, the cost of a single iteration is dominated by the extraction step, at $O(n^{2}+m)$. Therefore, the total worst-case complexity is $O(n^{3}+nm)$. Since $m=O(n^{2})$ for any simple graph, this simplifies to $O(n^{3})$. The initial sorting of vertices by non-increasing degree is a lower-order term that does not affect this bound.
We note that this is a theoretical worst-case analysis. In practice, EDQC is run as a time-bounded heuristic, and its empirical efficiency is demonstrated in our experimental section. ∎
4 Experimental Evaluation
We conducted a comprehensive experimental evaluation to validate the performance and robustness of EDQC. This section is organized as follows. We first detail the experimental setup, the baseline methods, and the datasets used. We then present the main results on effectiveness and efficiency, followed by a statistical analysis to confirm the significance of our findings. Subsequently, we provide an in-depth analysis of the algorithm, including an ablation study to verify the contribution of its key components and an empirical study to validate its core premise. The section concludes with a summary of our parameter sensitivity analysis.
4.1 Experimental Setup
All methods were implemented in C++ and compiled with g++ 9.4.0 using the ‘-O3’ optimization flag. The experiments were conducted on a machine equipped with an AMD EPYC 7H12 CPU (1.5 GHz base frequency) and 128 GB of RAM, running Ubuntu 20.04.
Each baseline was run with the parameter settings recommended in its original paper. For EDQC, we fixed the diffusion rounds at $T=2$ and the activation threshold at $\theta=10^{-4}$. These values were selected based on a sensitivity study, the details of which are provided in Section 4.8.
Due to inherent randomness, all non-deterministic methods (all except NBSim) were executed 10 times with random seeds from 1 to 10; NBSim, being deterministic, was run once. For the non-deterministic methods, we report their average quasi-clique size and standard deviation (STD) over the 10 runs, presented in the format ‘average size $\pm$ STD’, where the standard deviation is reported only if it is non-zero. All runs were subject to a 60-second timeout. If an algorithm failed to find any valid $\gamma$-quasi-clique in all 10 runs for an instance, we mark its result as ‘N/A’. The 60-second timeout reflects common practice in recent time-bounded heuristic studies lin2017reduction ; zheng2023farsighted .
4.2 Baseline Methods
We compared EDQC against four state-of-the-art algorithms:
  • NBSim/FastNBSim pang2024similarity : Similarity-based heuristics that expand subgraphs based on neighborhood overlap. FastNBSim accelerates NBSim via MinHash-based approximation.
  • NuQClqF1/NuQClqR1 liu2024optimization : Advanced variants of NuQClq chen2021nuqclq . These algorithms employ information feedback models that integrate historical solutions using fitness-weighted scores to guide local search.
4.3 Datasets
We evaluated all algorithms on 75 real-world datasets from SNAP snapnets and the Network Repository nr . The collection covers diverse domains, including web, social, and communication graphs. The datasets span a wide spectrum of scales, with vertex counts ($n$) ranging from 62 to over 16.7 million, and edge densities spanning from $8.41\times 10^{-2}$ down to as low as $1.87\times 10^{-7}$. This diversity provided a challenging testbed for evaluating the scalability and effectiveness of the algorithms.
4.4 Main Results
We now present the primary results comparing the effectiveness and efficiency of EDQC against the baselines. A key challenge in this comparison is that NBSim and FastNBSim do not accept a density threshold $\gamma$ as input. To facilitate a fair evaluation, we designed two separate experimental settings (Table 1 and Table 2). To save space, we omit instances where all four compared algorithms found solutions of the same average size.
Table 1: Quasi-clique sizes with the threshold $\gamma$ set by NBSim’s output.
Instance $\gamma$ NBSim NuQClqF1 NuQClqR1 EDQC
rt-retweet 1 3 4 4 4
web-polblogs 1 5 9 9 9
email-Eu-core 1 12 18 18 18
ego-facebook 0.99 71 92 92 92
ca-Erdos992 1 6 8 8 8
socfb-CMU 0.98 20 53 53 53
ech-WHOIS 0.98 57 79 79 79
web-indochina-2004 1 50 49 $\pm$ 3.16 50 50
socfb-UCSB37 0.99 40 64 64 64
socfb-UConn 0.99 22 58 58 58
socfb-UCLA 0.79 19 96 96 96
socfb-Berkeley13 1 21 42 42 42
socfb-Wisconsin87 0.98 12 46 46 46
soc-epinions 1 10 16 16 16
Cit-HepTh 1 12 22.9 $\pm$ 0.32 23 23
socfb-Indiana 0.85 27 94 94 94
socfb-UIllinois 1 14 57 57 57
Cit-HepPh 0.98 14 23 23 23
socfb-UF 0.99 24 67 66.8 $\pm$ 0.63 67
Email-Enron 0.97 10 25 25 25
socfb-Penn94 0.98 14 56 55.3 $\pm$ 2.21 56
soc-brightkite 0.99 36 45 45 45
socfb-OR 0.6 19 108 108 108
soc-slashdot 1 12 26 26 26
witter_combined 0.98 79 89.3 $\pm$ 11.7 89.3 $\pm$ 11.7 93
G_n_pin_pout 1 3 4 4 4
yahoo-msg 0.99 22 20.6 $\pm$ 2.07 20.6 $\pm$ 2.01 23
soc-douban 1 5 11 11 11
wave 1 5 6 6 6
Loc-Gowalla 0.99 31 23.8 $\pm$ 2.53 23 31
soc-gowalla 0.99 31 23 23.8 $\pm$ 2.53 31
web-Stanford 0.99 67 64.2 $\pm$ 3.43 59.9 $\pm$ 4.12 67.5 $\pm$ 0.71
cnr-2000 0.99 89 84.8 $\pm$ 3.61 81.8 $\pm$ 0.63 89
web-NotreDame 1 155 154.3 $\pm$ 0.48 154.2 $\pm$ 0.42 155
ca-MathSciNet 1 25 24.6 $\pm$ 0.84 24.7 $\pm$ 0.67 25
coPapersDBLP 1 337 333.5 $\pm$ 4.12 333.5 $\pm$ 4.12 337
auto 1 6 7 7 7
web-it-2004 1 432 431.8 $\pm$ 0.42 432 432
soc-delicious 0.58 22 80 80 80
ca-coauthors-dblp 1 337 333.5 $\pm$ 4.12 333.5 $\pm$ 4.12 337
eu-2005 0.99 405 191.3 $\pm$ 115.84 127.2 $\pm$ 26.92 405.1 $\pm$ 0.32
web-Google 0.99 48 39.2 $\pm$ 7.86 39.5 $\pm$ 7.41 48
rgg_n_2_20_s0 0.57 17 41 41 40.2 $\pm$ 0.79
rgg_n_2_21_s0 1 18 18.1 $\pm$ 0.57 18.1 $\pm$ 0.57 19
rgg_n_2_22_s0 1 19 18.8 $\pm$ 0.79 18.6 $\pm$ 0.7 20
rgg_n_2_23_s0 0.99 19 18.3 $\pm$ 0.48 19 $\pm$ 1.41 22
rgg_n_2_24_s0 0.98 20 19.7 $\pm$ 0.95 18.7 $\pm$ 0.48 23
Table 2: Quasi-clique sizes with $\gamma$ set by FastNBSim’s max density over 10 runs. For FastNBSim, stats are on successful runs only (rate shown in parentheses).
Instance $\gamma$ FastNBSim NuQClqF1 NuQClqR1 EDQC
soc-dolphins 0.83 4 (4/10) 6 6 6
rt-retweet 1 3 (7/10) 4 4 4
ca-netscience 1 8.1 $\pm$ 0.88 (10/10) 9 9 9
web-polblogs 0.93 6 (2/10) 13 13 13
rt-twitter-copen 1 3.6 $\pm$ 0.53 (7/10) 4 4 4
email-Eu-core 0.94 21 (1/10) 29 29 29
web-edu 1 27.4 $\pm$ 3.1 (10/10) 30 30 30
ego-facebook 0.96 102 (1/10) 123 123 123
ca-GrQc 1 37.8 $\pm$ 3.2 (4/10) 44 44 44
web-spam 1 15.5 $\pm$ 0.71 (2/10) 20 20 20
ca-Erdos992 1 5 (6/10) 8 8 8
socfb-CMU 0.97 28 (1/10) 56 56 56
ech-WHOIS 0.96 73 (1/10) 91 91 91
ca-HepPh 1 202.7 $\pm$ 24.04 (10/10) 239 239 239
web-indochina-2004 1 49.9 $\pm$ 0.32 (10/10) 49 $\pm$ 3.16 50 50
web-BerkStan 1 26.7 $\pm$ 2.83 (10/10) 29 29 29
socfb-UCSB37 0.98 47 (1/10) 70 70 70
web-webbase-2001 1 32.9 $\pm$ 0.32 (10/10) 33 33 33
socfb-UConn 0.97 34 (1/10) 65 65 65
ca-AstroPh 1 51.8 $\pm$ 3.12 (10/10) 57 57 57
socfb-UCLA 0.98 49 (1/10) 63 63 63
ca-CondMat 1 22.63 $\pm$ 2.83 (8/10) 26 26 26
socfb-Berkeley13 0.95 29 (1/10) 62 62 62
socfb-Wisconsin87 0.91 34 (1/10) 65 65 65
soc-epinions 0.95 16 (1/10) 21 21 21
Cit-HepTh 0.92 26 (1/10) 41 41 41
socfb-Indiana 0.95 40 (1/10) 71 70.8 $\pm$ 0.63 71
socfb-UIllinois 0.94 48 (1/10) 92 92 92
Cit-HepPh 0.98 20 (1/10) 23 23 23
socfb-UF 0.94 69 (1/10) 93 93 93
Email-Enron 0.96 13 (1/10) 26 26 26
socfb-Penn94 0.94 42 (1/10) 63.4 $\pm$ 3.37 63.3 $\pm$ 3.59 65
soc-brightkite 0.97 43 (1/10) 53 53 53
socfb-OR 0.92 24 (1/10) 50.9 $\pm$ 0.32 50.9 $\pm$ 0.32 51
soc-slashdot 0.97 10 (1/10) 35 35 35
wordnet-words 1 29.2 $\pm$ 2.71 (6/10) 32 32 32
witter_combined 0.97 87 (1/10) 99 99 99
G_n_pin_pout 0.5 4 (3/10) 7.9 $\pm$ 0.32 7.9 $\pm$ 0.32 8
yahoo-msg 0.96 24 (1/10) 23.4 $\pm$ 1.84 23.9 $\pm$ 1.1 25
web-sk-2005 0.99 83 (3/10) 84 84 84
soc-douban 0.72 9 (1/10) 23 23 23
wave 0.71 8 (3/10) 14 14 14
web-arabic-2005 1 90.7 $\pm$ 5.36 (10/10) 102 102 102
Loc-Gowalla 1 23.5 $\pm$ 0.71 (2/10) 21 21 29
soc-gowalla 0.99 25 (1/10) 23 23.8 $\pm$ 2.53 31
ca-dblp-2010 1 70.6 $\pm$ 4.86 (10/10) 75 75 75
ca-citeseer 1 81.6 $\pm$ 5.36 (10/10) 87 87 87
coAuthorsCiteseer 1 81.6 $\pm$ 5.36 (10/10) 87 87 87
web-Stanford 0.99 62 (1/10) 64.2 $\pm$ 3.43 59.9 $\pm$ 4.12 67.5 $\pm$ 0.71
coAuthorsDBLP 1 103 $\pm$ 9.7 (10/10) 115 115 115
ca-dblp-2012 1 104 $\pm$ 9.52 (10/10) 114 114 114
com-dblp 1 101.4 $\pm$ 5.76 (10/10) 114 114 114
cnr-2000 1 84 (1/10) 81.9 $\pm$ 1.45 81.8 $\pm$ 1.55 84
web-NotreDame 1 153.5 $\pm$ 1.43 (10/10) 154.3 $\pm$ 0.48 154.2 $\pm$ 0.42 155
ca-MathSciNet 1 23.8 $\pm$ 0.79 (10/10) 24.6 $\pm$ 0.84 24.5 $\pm$ 0.85 25
coPapersDBLP 1 316.9 $\pm$ 23 (10/10) 333 $\pm$ 3.94 333 $\pm$ 3.94 337
auto 0.77 9.5 $\pm$ 0.71 (2/10) 11 11 11
web-it-2004 1 431.7 $\pm$ 0.48 (10/10) 431.8 $\pm$ 0.42 432 432
soc-delicious 0.88 30 (1/10) 31.4 $\pm$ 3.86 31.4 $\pm$ 3.86 37
ca-coauthors-dblp 1 316.9 $\pm$ 23 (10/10) 333.5 $\pm$ 4.12 330.2 $\pm$ 10.01 337
soc-digg 0.9 76 (1/10) 104 $\pm$ 13.7 98.8 $\pm$ 12.56 116.9 $\pm$ 0.32
eu-2005 0.99 387 (1/10) 191.3 $\pm$ 115.84 127.2 $\pm$ 26.92 405.1 $\pm$ 0.32
web-Google 1 40.7 $\pm$ 3.21 (3/10) 33.7 $\pm$ 2.63 37.1 $\pm$ 6.08 44
soc-pokec 0.99 31 (1/10) 24.6 $\pm$ 2.37 23.8 $\pm$ 1.55 25
rgg_n_2_22_s0 0.98 19 (1/10) 20.1 $\pm$ 0.99 20 $\pm$ 0.94 21
rgg_n_2_23_s0 0.99 21 (1/10) 18.3 $\pm$ 0.48 19 $\pm$ 1.41 22
hugetrace-00010 0.5 4 (10/10) N/A N/A 5
hugetrace-00020 0.5 4 (10/10) N/A N/A 5
rgg_n_2_24_s0 0.95 20 (1/10) 20.6 $\pm$ 1.26 20.9 $\pm$ 1.37 25
Effectiveness and Stability.
The results in Table 1 and Table 2 show the superior effectiveness and stability of EDQC. A quantitative summary highlights this advantage. In the first setting (Table 1), across the 47 instances where results differed, EDQC achieves or ties for the best average size in 46 instances. This compares favorably to the other methods, where NuQClqR1 achieved or tied for the best size in 29 instances, followed by NuQClqF1 (28 instances) and NBSim (10 instances). Similarly, in the second setting (Table 2), EDQC achieves or ties for the best size in 68 out of 69 instances. For comparison, the counts for the other algorithms were: NuQClqR1 (46 instances), NuQClqF1 (45 instances), and FastNBSim (2 instances).
Beyond average size, EDQC demonstrates exceptional stability. It consistently reports a standard deviation close to zero (e.g., $0.32$ on the large eu-2005 graph), even on complex instances. This is in stark contrast to the NuQClq variants, which frequently exhibit high variance on the same instances (e.g., a STD of $115.84$ for NuQClqF1 on eu-2005). Furthermore, the results highlight the unreliability of the similarity-based methods. FastNBSim, for instance, often fails to achieve its own best-found density (succeeding only 1/10 times on many graphs). This lack of explicit density control makes them less suitable for applications requiring highly cohesive structures, a task where EDQC’s reliability is a distinct advantage.
Figure 2: Runtime comparison of EDQC against baselines on a log-log scale.
Efficiency.
The efficiency of EDQC is compared against baselines in Figure 2, where, following prior work sunnumds , instances where the runtime of both algorithms is below 0.1 seconds are excluded. On this log-log plot, points falling below the $1\times$, $10\times$, and $100\times$ lines indicate that EDQC is faster than the baseline by the corresponding factor or more. The results clearly show that EDQC is substantially more efficient than the high-performing local search methods, NuQClqF1 and NuQClqR1, with a large number of points falling below the $10\times$ line, signifying a speedup of over an order of magnitude. In comparison to the fast similarity-based heuristics (NBSim and FastNBSim), EDQC’s runtime remains highly competitive, as most of their corresponding points cluster around the $1\times$ line. Overall, this demonstrates that EDQC strikes an excellent balance between achieving state-of-the-art solution quality and maintaining high computational efficiency.
4.5 Statistical Analysis
(a) Comparison in the NBSim setting.
(b) Comparison in the FastNBSim setting.
Figure 3: Critical Difference diagrams for algorithm ranks on all 75 datasets (significance level 0.05). A horizontal bar connects algorithms with no statistically significant difference.
To assess statistical significance, we applied the Friedman test friedman1937use on the algorithm rankings across all 75 datasets. The null hypothesis that all algorithms perform equally was rejected in both settings ($p<0.05$), allowing us to proceed with the post-hoc Nemenyi test demvsar2006statistical to identify pairwise performance differences.
The results are visualized using Critical Difference (CD) diagrams in Figure 3. In these diagrams, algorithms are plotted on an axis according to their average rank across all datasets; a lower rank (further to the left) is better. Groups of algorithms whose performance is not statistically significantly different are connected by a horizontal bar, the length of which corresponds to the critical difference.
The top plot in Figure 3 shows the comparison for the NBSim setting. EDQC achieves the best average rank, and this leading position is statistically significant. The post-hoc test reveals that EDQC’s performance is significantly better than all three other baselines: NuQClqF1, NuQClqR1, and NBSim. A similar trend is observed in the FastNBSim setting, shown in the bottom plot. Once again, EDQC obtains the best average rank. Its superiority is also confirmed to be statistically significant, as it significantly outperforms NuQClqF1, NuQClqR1, and FastNBSim.
In summary, the statistical tests provide conclusive evidence for the superiority of EDQC. In both experimental settings, EDQC not only achieves the best average rank but also demonstrates a statistically significant performance advantage over all compared state-of-the-art baselines. These results provide strong statistical evidence that EDQC represents a reliably superior approach for this task.
4.6 Analysis of EDQC Components
To isolate and verify the contribution of each key component, we conducted an ablation study comparing the full EDQC against five variants across 112 unique experimental settings (consolidating settings where both configurations yielded the same $\gamma$ value). These variants were designed to test the core aspects of our algorithm: the subgraph identification strategy (EDQC-1), the energy-guided ranking (EDQC-2, EDQC-3), the refinement phase (EDQC-4), and the nature of the diffusion process itself (EDQC-5).
The results, summarized in Table 3, provide conclusive evidence that each component is essential for EDQC’s superior performance. The most dramatic performance degradation occurred with EDQC-5, which replaced our adaptive diffusion with a naive random process, and EDQC-1, which substituted conductance-based identification with a max-density heuristic. The full EDQC algorithm proved superior in 106 and 69 instances, respectively. These results underscore that our principled diffusion and extraction mechanisms form the foundational pillars of the algorithm’s success. The importance of the energy-guided ranking was also clearly demonstrated; disrupting the descending energy order with a random (EDQC-2) or ascending (EDQC-3) ranking severely hampered performance, leading to 56 and 70 wins for the full EDQC. Finally, disabling the refinement phase (EDQC-4) also resulted in a noticeable performance drop in 16 instances, confirming that this final step is vital for optimizing the solution size.
Additionally, we confirmed the robustness of our synchronous design. As variants with a fixed processing order for active vertices produced identical results to the full EDQC, the order-independence guaranteed by the synchronous design was empirically verified.
Table 3: Comparison of the full EDQC with its five simplified variants across 112 experimental settings. #bet and #wor represent the number of instances where EDQC achieves better and worse quasi-clique sizes, respectively.
Variant Description #bet #wor
(EDQC ¿ Variant) (EDQC ¡ Variant)
EDQC-1 No Conductance 69 0
EDQC-2 Random Ranking 56 1
EDQC-3 Ascending Ranking 70 1
EDQC-4 No Refinement Phase 16 1
EDQC-5 Naive Random Diffusion 106 0
Figure 4: Subgraph density vs. total retained energy on eight representative datasets. Each blue dot represents one of 1,000 randomly sampled subgraphs of a fixed size; the red marks the quasi-clique of the same size found by EDQC.
4.7 Empirical Validation of the Energy-Density Premise
To empirically validate our algorithm’s core premise that energy concentrates in dense regions, we analyzed the relationship between subgraph density and energy retention on eight representative datasets. For each dataset, we compared the quasi-clique found by EDQC against 1,000 randomly sampled subgraphs of the same size. The results, visualized in Figure 4, reveal a strong positive correlation between density and retained energy, confirmed by high Pearson correlation coefficients (typically >0.90>0.90). More importantly, the quasi-clique discovered by EDQC (red marks) consistently appears as a top-right outlier, indicating that it captures a disproportionately high amount of energy even among other subgraphs of similar density. This strong alignment provides the empirical justification for using the energy ranking as a powerful guide in the discovery of quasi-cliques.
4.8 Parameter Sensitivity
This section provides the detailed results of our parameter sensitivity analysis, which was designed to evaluate the impact of the number of diffusion rounds $T$ and the activation threshold $\theta$ on EDQC’s performance. For this study, we selected three representative datasets to cover different scales: ego-facebook (a small graph, $|V|\leq 10^{4}$), rgg_n_2_19_s0 (a medium graph, $10^{4}<|V|\leq 10^{6}$), and web-Google (a large graph, $|V|>10^{6}$).
On each of these graphs, we evaluated EDQC by testing 15 combinations of $T\in\{1,2,3\}$ and $\theta\in\{10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3},10^{-2}\}$. This evaluation was repeated under three different density thresholds, $\gamma\in\{0.7,0.8,0.9\}$, to ensure a comprehensive analysis. We report both the discovered quasi-clique size and the runtime for each configuration. The results for the small, medium, and large graphs are presented in Figure 57.
Figure 5: Parameter sensitivity analysis for EDQC on the small representative graph (ego-facebook). Each of the three rows corresponds to a different density threshold $\gamma$. For each $\gamma$, the left plot shows the resulting quasi-clique size versus the activation threshold $\theta$ (on a log scale), and the right plot shows the corresponding runtime.
Figure 6: Parameter sensitivity analysis on the medium representative graph (rgg_n_2_19_s0). The layout of the plots is identical to that described in the caption of Figure 5.
Figure 7: Parameter sensitivity analysis on the large representative graph (web-Google). The layout of the plots is identical to that described in the caption of Figure 5.
Across all three datasets, a consistent pattern emerges. In terms of solution quality (size), configurations with $T=2$ and $T=3$ consistently and significantly outperform $T=1$, especially for smaller values of the activation threshold ($\theta\leq 10^{-3}$). As $\theta$ increases, making vertex activation stricter, the performance of all settings tends to decline. Regarding efficiency, the runtime plots clearly show that the computational cost increases with $T$. While $T=3$ occasionally finds a marginally larger quasi-clique than $T=2$, its runtime is substantially higher across all settings. For instance, in the large graph setting (Figure 7), the runtime for $T=3$ can be several times that of $T=2$ for little to no improvement in solution size.
Based on this analysis, we conclude that the configuration $T=2$ provides the best trade-off between solution quality and computational cost. Paired with a small activation threshold such as $\theta=10^{-4}$, which consistently lies in the stable, high-performance region shown in the plots, this configuration represents an effective and efficient choice. This justifies our use of $T=2$ and $\theta=10^{-4}$ for all main experiments presented in the paper.
5 Conclusion
In this paper, we introduced EDQC, a novel algorithm for quasi-clique discovery designed to mitigate key limitations of existing methods, such as their sensitivity to random seeds and lack of explicit edge-density guarantees. We demonstrated that the effectiveness of EDQC stems from its two-stage design. First, an adaptive energy diffusion process generates an energy ranking that highlights structurally cohesive regions. Then, guided by this ranking, a principled extraction mechanism based on conductance minimization identifies and refines a high-quality subgraph. To the best of our knowledge, EDQC is the first method to incorporate energy diffusion into this task, offering a robust alternative to traditional heuristics.
Our extensive experiments on 75 real-world graphs confirmed the effectiveness of this approach. The results demonstrated that EDQC finds larger quasi-cliques on most datasets, with consistently lower variance across runs, highlighting its superior stability. Furthermore, statistical tests confirmed the significance of these findings, and our analysis showed that EDQC maintains a competitive runtime, striking an excellent balance between solution quality and efficiency.
Future work can extend the EDQC framework in several promising directions. Generalizing the diffusion model for weighted and attributed graphs is an important next step. Applying EDQC to downstream tasks, such as anomaly detection in financial networks and community detection in social graphs, could further validate its practical utility. Finally, exploring more advanced refinement strategies may lead to further improvements in solution quality.

### v3

@@
-Quasi-Clique Discovery via Energy Diffusion
-Abstract
-Discovering quasi-cliques—subgraphs whose edge density exceeds a given threshold—is a fundamental task in graph mining with applications to web spam detection, fraud screening, and e-commerce recommendation.
-However, existing methods for quasi-clique discovery on large-scale web graphs are often sensitive to random seeds or lack of explicit edge-density guarantees, making the task challenging in practice.
-This paper presents EDQC, an energy diffusion–based method for quasi-clique discovery.
-EDQC first employs an adaptive energy diffusion process to generate an energy ranking that highlights structurally cohesive regions.
-Guided by this energy ranking, the algorithm identifies a high-quality subgraph by minimizing conductance, a standard measure from community detection.
-This subgraph is then refined to meet the specified density threshold.
-Extensive experiments on 75 real-world graphs show that EDQC finds larger quasi-cliques on most datasets, with consistently lower variance across runs and competitive runtime.
-To the best of our knowledge, EDQC is the first method to incorporate energy diffusion into quasi-clique discovery.
-1 Introduction
-Dense subgraph models are fundamental to graph mining on large-scale web graphs, enabling web spam detection gibson2005discovering ; lin2021mining , fraud screening zheng2019author ; fang2022densest , and e-commerce recommendation schafer2001commerce ; song2021subgraph .
-Among these models, the clique represents the most stringent form, requiring that every pair of vertices be directly connected to each other saito2007large .
-However, due to noise and missing links in real-world graphs, exact cliques are rarely observed.
-To relax this strict requirement, several approximate dense subgraph models have been proposed, including quasi-cliques abello2002massive ; pei2005mining , $k$-clubs mokken1979cliques , and $k$-plexes seidman1978graph .
-In particular, the $\gamma$-quasi-clique, defined by a minimum edge-density threshold, is one of the most widely adopted formulations chen2021nuqclq ; pang2024similarity ; liu2024optimization .
-Given an undirected graph $G$ and an edge-density threshold $\gamma\in(0,1]$, the Maximum Quasi-Clique Problem (MQCP) seeks the largest vertex set $S$ such that the induced subgraph $G[S]$ exhibits an edge density of at least $\gamma$.
-This edge-density criterion provides a more direct measure of structural cohesion compared to the degree-based definitions utilized by methods such as FastQC konar2024optimal and IterQC liu2008effective .
-However, the MQCP is NP-hard even for fixed values of $\gamma$ pattillo2013maximum and does not possess the hereditary property that allows for efficient pruning in problems related to cliques xia2025maximum .
-Nonetheless, quasi-clique discovery remains highly valuable, as real-world graphs typically exhibit dense yet incomplete connectivity rather than perfect cliques, making quasi-cliques crucial for identifying significant or anomalous substructures.
-Figure 1: A clique and a 0.8-quasi-clique.
-One practical application of quasi-clique discovery is identifying link farm spam pages in web graphs pang2024similarity , where densely connected spam pages form $\gamma$-quasi-cliques, enabling the identification of suspicious link farms even under noise and missing links.
-As shown in Figure 1, the dashed edges in the left subgraph form a size-4 clique, representing a fully connected link farm.
-In the right subgraph, the dashed edges form a 0.8-quasi-clique, still suspicious despite one missing link, capturing a potential link farm under partial observation or evasion.
-Despite recent progress, discovering quasi-cliques in real large-scale web graphs remains challenging due to both scale and combinatorial complexity.
-While exact algorithms for MQCP, typically based on integer programming or techniques such as branch-and-bound uno2010efficient ; pastukhov2018maximum ; ribeiro2019exact ; rahman2024fast , guarantee optimal solutions, their exponential time complexity restricts their applicability to large graphs.
-As a practical alternative, numerous heuristic methods have been proposed to enable scalable quasi-clique discovery, which can be roughly grouped into four categories.
-First, greedy expansion techniques, such as the greedy strategy with local refinement introduced in tsourakakis2013denser and the restart-enabled search in RIG oliveira2013construction , prioritize speed but are vulnerable to becoming trapped in local optima.
-Second, metaheuristic frameworks like TSQC djeddi2019extension and OBMA zhou2020opposition employ tabu search and opposition-based learning, offering enhanced robustness at a higher computational cost.
-Third, local search algorithms, including NuQClq chen2021nuqclq and its variants NuQClqF1 and NuQClqR1 liu2024optimization , utilize bounded configuration checking and adaptive rules for candidate refinement.
-However, despite their sophisticated local optimizations, these methods often exhibit instability due to their dependence on random seeds.
-Lastly, similarity-based heuristics such as NBSim and its accelerated variant, FastNBSim, utilize vertex neighborhood overlap to identify and expand quasi-cliques pang2024similarity .
-While both methods are highly efficient, FastNBSim introduces MinHash-based approximations, making it sensitive to random seeds and potentially less stable.
-Moreover, neither method supports a user-specified edge-density threshold.
-Instead, the edge-density of the returned subgraph emerges implicitly from similarity parameters.
-Consequently, there is no guarantee that the output meets a specified minimum edge-density requirement.
-This limits their suitability for applications that require strict control over quasi-clique edge-density thresholds.
-Among these four categories, local search and similarity-based methods have been the most competitive for discovering large quasi-cliques.
-The state-of-the-art algorithms for MQCP include NBSim and its accelerated variant, FastNBSim pang2024similarity , as well as the local search algorithms NuQClqF1 and NuQClqR1 liu2024optimization .
-However, these methods remain sensitive to random seeds or lack of explicit edge-density guarantees.
-Energy diffusion is a natural candidate for addressing the challenges above.
-Diffusion methods such as PageRank page1999pagerank and the heat kernel chung2007heat have shown strong performance on structure-aware tasks, e.g., vertex ranking and community detection kloster2014heat ; gleich2015multilinear , where diffused mass tends to localize within cohesive regions.
-These tasks overlap with quasi-clique discovery in highlighting tightly interlinked vertices, yet they are not equivalent because quasi-cliques must satisfy a specified minimum edge-density on induced subgraphs.
-This connection motivates exploring energy diffusion for quasi-clique discovery to reduce sensitivity to random seeds in practice and to furnish a stable candidate region for density-constrained extraction.
-This paper introduces EDQC, an energy diffusion-based method for quasi-clique discovery.
-EDQC begins with an adaptive energy diffusion phase, where the energy transfer rate is dynamically adjusted based on the local energy distribution.
-This phase runs for a fixed number of iterations from each source vertex, generating an energy ranking that highlights structurally cohesive regions.
-Guided by this energy ranking, EDQC then employs a principled extraction mechanism from community detection theory.
-It applies a sweep cut to the energy-ranked vertices to identify a high-quality subgraph that minimizes conductance.
-This subgraph is subsequently refined through a structure-driven procedure to guarantee it meets the specified density threshold.
-Extensive experiments on 75 real-world graphs, including classic web graphs, demonstrate that EDQC discovers larger quasi-cliques on most datasets, with consistently lower variance and competitive runtime.
-To the best of our knowledge, EDQC is the first method to incorporate energy diffusion into quasi-clique discovery.
-2 Preliminaries
-All graphs considered in this paper are undirected and simple (no loops or parallel edges).
-Let $G=(V,E)$ represent a graph with a vertex set $V$ and an edge set $E$.
-Two vertices $u$ and $v$ within $V$ are termed adjacent if $uv\in E$, and one vertex is referred to as a neighbor of the other.
-The set of all neighbors for a vertex $u$ is known as its neighborhood, denoted $N(u)$, while the closed neighborhood of $u$ is denoted $N[u]=N(u)\cup\{u\}$.
-The degree of a vertex $u$, denoted as $d(u)$, indicates the number of its neighbors, i.e., $d(u)=|N(u)|$.
-Let $n=|V|$ and $m=|E|$.
-For a vertex set $S\subseteq V$, the induced subgraph is $G[S]$ with vertex set $S$ and edge set $E(S)=\{\{u,v\}\in E\mid u,v\in S\}$.
-Let $m(S)=|E(S)|$ denote the number of internal edges in $G[S]$, and define the internal degree of $u\in S$ by
-$d_{S}(u)=|N(u)\cap S|.$
-The edge density of $S$ (and of $G[S]$) is
-$\eta(S)=\begin{cases}\dfrac{2\,m(S)}{|S|(|S|-1)},&|S|\geq 2,\\[6.0pt] 0,&otherwise,\end{cases}$
-and $\eta(G)$ is shorthand for $\eta(V)$. Throughout this paper, unless otherwise stated, density means edge density.
+Cohesive Group Discovery in Interaction Graphs under Explicit Density Constraints
+Abstract.
+Discovering cohesive groups is a fundamental primitive in graph-based recommender systems, underpinning tasks such as social recommendation, bundle discovery, and community-aware modeling.
+In interaction graphs, cohesion is often modeled as the $\gamma$-quasi-clique, an induced subgraph whose internal edge density meets a user-defined threshold $\gamma$.
+This formulation provides explicit control over within-group connectivity while accommodating the sparsity inherent in real-world data.
+This paper presents EDQC, an effective framework for cohesive group discovery under explicit density constraints.
+EDQC leverages a lightweight energy diffusion process to rank vertices for localizing promising candidate regions.
+Guided by this ranking, the framework extracts and refines a candidate subgraph to ensure the output strictly satisfies the target density requirement.
+Extensive experiments on 75 real-world graphs across varying density thresholds demonstrate that EDQC identifies the largest mean $\gamma$-quasi-cliques in the vast majority of cases, achieving lower variance than the state-of-the-art methods while maintaining competitive runtime.
+Furthermore, statistical analysis confirms that EDQC significantly outperforms the baselines, underscoring its robustness and practical utility for cohesive group discovery in graph-based recommender systems.
+Cohesive group discovery, quasi-clique, graph-based recommender systems
+copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ACM SIGIR Conference 2026; July 20–24, 2026; Melbourne — Naarm, Australiaisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
+1. Introduction
+Cohesive group discovery plays a fundamental role in graph-based recommender systems, enabling tasks such as social recommendation (Alina Christensen and Schiaffino, 2014), bundle discovery (Nguyen et al., 2024), and community-aware modeling (Guan et al., 2021).
+In many recommender settings, user–item interactions or user–user relations are naturally modeled as interaction graphs, where vertices denote users or items and edges capture affinity signals such as co-consumption or behavioral similarity (Schafer et al., 2001; Zhou, 2025).
+This representation encompasses various network types, including social, communication, citation, collaboration, and web graphs.
+Within such graphs, cohesive groups are often instantiated as dense induced subgraphs, supporting candidate expansion, group-aware ranking, and coordinated abnormal behavior detection.
+Among dense subgraph models, the clique is the most stringent, requiring every pair of vertices to be connected (Saito et al., 2007).
+However, in real-world networks, noise, missing edges, and incomplete connectivity make exact cliques rare.
+To address this, several relaxed models have been proposed, including quasi-cliques (Abello et al., 2002; Pei et al., 2005), $k$-clubs (Mokken et al., 1979), and $k$-plexes (Seidman and Foster, 1978).
+In particular, the edge-density-based formulation, often referred to as the $\gamma$-quasi-clique, is widely adopted for its intuitive and explicit notion of structural cohesion (Chen et al., 2021; Pang et al., 2024; Liu et al., 2024).
+Given an undirected graph $G$ and an edge-density threshold $\gamma\in(0,1]$, the Maximum Quasi-Clique Problem (MQCP) asks for a largest vertex set $S$ such that the induced subgraph $G[S]$ has edge density at least $\gamma$.
+Throughout the paper, density refers to edge density, and quasi-clique refers to the $\gamma$-quasi-clique unless otherwise stated.
+Unlike degree-based criteria used in FastQC (Konar and Sidiropoulos, 2024) and IterQC (Liu and Wong, 2008), MQCP enforces an explicit, user-specified density threshold.
+It is also distinct from community detection, which typically optimizes a global objective (e.g., modularity or conductance) to partition or cover the entire graph.
+MQCP instead extracts a single dense local vertex set, aligning with settings where one seeks a single cohesive set of users or items rather than a full partition.
+In recommender applications, $\gamma$ serves as a minimum cohesion requirement that explicitly controls within-group connectivity.
+Hence, methods that do not enforce a user-specified $\gamma$ may produce subgraphs that fail to satisfy the target density, limiting their utility as reliable constraint-aware components for recommendation tasks.
+MQCP is NP-hard even for a fixed $\gamma$ (Pattillo et al., 2013) and lacks the hereditary property used for pruning in clique-finding algorithms (Xia et al., 2025).
+In particular, a subset of a $\gamma$-quasi-clique is not necessarily a $\gamma$-quasi-clique.
+Despite these challenges, MQCP remains practically relevant, since $\gamma$-quasi-cliques capture compact interest groups and suspicious behavioral clusters in real-world graphs with dense but imperfect connectivity.
+Figure 1 illustrates a typical situation in interaction graphs.
+In the left subgraph, the dashed edges form a size-4 clique, whereas in the right subgraph one edge is missing but the dashed edges still constitute a $0.8$-quasi-clique.
+In practice, interactions are often noisy or partially observed, so nearly-complete groups like this should still be treated as coherent interest groups.
+This motivates studying $\gamma$-quasi-cliques as an explicit density-constrained model of structural cohesion.
+Figure 1. A size-4 clique and a 0.8-quasi-clique. Dashed edges indicate observed interaction affinities.
+Despite recent progress, discovering quasi-cliques in large graphs remains challenging due to both scale and combinatorial complexity.
+Exact MQCP solvers based on integer programming and branch-and-bound (Uno, 2010; Pastukhov et al., 2018; Ribeiro and Riveaux, 2019; Rahman et al., 2024) provide optimality guarantees but are restricted to small graphs in practice.
+Consequently, heuristics are often preferred for real-world graphs, and existing methods can be roughly grouped into four categories.
+First, greedy expansion methods grow candidates with lightweight refinement (Tsourakakis et al., 2013; Oliveira et al., 2013).
+While fast, they are prone to getting trapped in poor local optima.
+Second, metaheuristic frameworks such as TSQC and OBMA (Djeddi et al., 2019; Zhou et al., 2020) enhance exploration using strategies like tabu search or opposition-based learning, but typically incur higher computational overhead, which limits scalability on large graphs.
+Third, local search algorithms, notably NuQClq and its variants (Chen et al., 2021; Liu et al., 2024), iteratively refine candidates with adaptive move rules (e.g., bounded configuration checking); however, their performance can vary substantially across random seeds on large-scale instances.
+This seed sensitivity hinders reproducibility and yields unstable outputs across runs, which is undesirable in recommender systems.
+Fourth, similarity-based heuristics such as NBSim and FastNBSim (Pang et al., 2024) exploit neighborhood overlap for efficiency, but they do not take a user-specified edge-density threshold $\gamma$ as input.
+Instead, density is controlled indirectly via similarity parameters, so the output is not guaranteed to satisfy a target $\gamma$.
+As a result, their outputs may be dense under an internal similarity criterion yet still violate a required edge-density threshold, making them unsuitable when $\gamma$ is a hard constraint.
+Among these categories, local search and similarity-based heuristics are often among the strongest baselines for mining large quasi-cliques.
+In particular, representative methods include NBSim/FastNBSim (Pang et al., 2024) and the local search algorithms NuQClqF1/NuQClqR1 (Liu et al., 2024).
+However, these methods either exhibit sensitivity to random seeds, or do not take a user-specified density threshold $\gamma$ as an explicit constraint and thus cannot guarantee meeting a target $\gamma$.
+Energy diffusion provides a natural mechanism to improve the stability of quasi-clique discovery.
+Diffusion-based techniques such as PageRank (Page et al., 1999) and the heat kernel (Chung, 2007) have been effective for structure-aware tasks, including vertex ranking and local community detection (Kloster and Gleich, 2014; Gleich et al., 2015).
+A simple intuition is that short-range propagation tends to retain more mass within well-connected regions due to repeated exchanges along internal edges, which can be used to produce a locality-aware ranking from a source vertex.
+However, classical diffusion models are designed to approximate a probability distribution and are not tailored to extracting a vertex set that explicitly satisfies a density threshold $\gamma$.
+To bridge this gap, we present EDQC, an effective framework that leverages energy diffusion for quasi-clique discovery.
+By utilizing lightweight primitives and sidestepping the iterative overhead often associated with metaheuristic search, EDQC is designed to be practical across different graph sizes.
+EDQC first runs an adaptive energy diffusion process to obtain an energy ranking, where the release rate of each vertex is adjusted by the energy in its neighborhood to concentrate scores within locally cohesive regions.
+While the ranking highlights promising vertices, it does not define a subgraph boundary.
+We therefore apply a conductance-based sweep cut to extract a well-separated candidate region, followed by a refinement step that prunes low-degree vertices and greedily adds qualifying vertices.
+The refinement enforces feasibility, ensuring that the output meets the user-specified density threshold $\gamma$.
+Our main contributions are summarized as follows:
+  • To our knowledge, we are the first to integrate energy diffusion into a heuristic framework for MQCP, using adaptive diffusion-based scoring to localize cohesive regions in interaction graphs at scale.
+  • We design a diffusion-guided framework that converts local rankings into feasible $\gamma$-quasi-cliques through conductance-based extraction and density-constrained refinement, ensuring that the output satisfies the user-specified density threshold $\gamma$.
+  • We evaluate EDQC on 75 real-world graphs with ablation studies and statistical significance tests, showing that it finds larger quasi-cliques than strong baselines on most datasets, with lower variance across runs and competitive runtime.
+2. Preliminaries
+All graphs are undirected and simple (no loops or parallel edges).
+Let $G=(V,E)$ denote a graph with vertex set $V$ and edge set $E$, where $n=|V|$ and $m=|E|$.
+For a vertex $u\in V$, let $N(u)$ be its neighborhood, $N[u]=N(u)\cup\{u\}$ its closed neighborhood, and $d(u)=|N(u)|$ its degree.
+For a set $S\subseteq V$, its volume is $\mathrm{vol}(S)=\sum_{u\in S}d(u)$.
+Let $G[S]$ be the induced subgraph with vertex set $S$ and edge set $E(S)=\{\{u,v\}\in E\mid u,v\in S\}$.
+Let $m(S)=|E(S)|$ be the number of internal edges in $G[S]$, and for $u\in S$ define the internal degree $d_{S}(u)=|N(u)\cap S|$.
+The edge density of $G[S]$ is
+$\eta(G[S])=\begin{cases}\dfrac{2m(S)}{|S|(|S|-1)},&|S|\geq 2,\\ 0,&\text{otherwise}.\end{cases}$
 Definition 0 ($\gamma$-Quasi-Clique).
-Given $G=(V,E)$ and a density threshold $\gamma\in(0,1]$. A vertex set $S\subseteq V$ induces a $\gamma$-quasi-clique if the induced subgraph $G[S]$ has edge density at least $\gamma$, i.e.,
-$\eta(G[S])\;\geq\;\gamma.$
-For brevity, we also refer to $G[S]$ itself as a $\gamma$-quasi-clique.
-A $\gamma$-quasi-clique $G[S]$ is maximal if there is no $v\in V\setminus S$ such that $\eta\!\left(G[S\cup\{v\}]\right)\geq\gamma$.
+Given $G=(V,E)$ and a density threshold $\gamma\in(0,1]$, a vertex set $S\subseteq V$ induces a $\gamma$-quasi-clique if the induced subgraph $G[S]$ has edge density at least $\gamma$, i.e., $\eta(G[S])\;\geq\;\gamma.$ For brevity, we also refer to $G[S]$ itself as a $\gamma$-quasi-clique.
 Definition 0 (Maximum Quasi-Clique Problem).
@@
 The induced subgraph $G[S^{*}]$ is a maximum $\gamma$-quasi-clique.
-When $\gamma=1$, MQCP reduces to the classical Maximum Clique Problem.
-For any fixed $\gamma$, MQCP is NP-hard and may admit multiple optimal solutions pattillo2013maximum .
+When $\gamma=1$, MQCP reduces to the classical Maximum Clique Problem, and for any fixed $\gamma$ the problem is NP-hard and may admit multiple optimal solutions (Pattillo et al., 2013).
 Definition 0 (Conductance).
-Conductance is a standard measure used to evaluate how well a vertex set forms a community andersen2006local .
-For a non-empty, proper subset of vertices $S\subset V$, its volume is defined as $\mathrm{vol}(S)=\sum_{u\in S}d(u)$, and the cut between $S$ and its complement is $\mathrm{cut}(S,V\setminus S)=|\{\{u,v\}\in E\mid u\in S,v\in V\setminus S\}|$.
+Conductance measures how well a vertex set is separated from the rest of the graph and is widely used in local clustering (Andersen et al., 2006).
+For a non-empty, proper subset of vertices $S\subset V$, the cut between $S$ and its complement is defined as $\mathrm{cut}(S,V\setminus S)=|\{\{u,v\}\in E\mid u\in S,v\in V\setminus S\}|$.
 The conductance of $S$ is then:
 $\Phi(S)=\frac{\mathrm{cut}(S,V\setminus S)}{\min(\mathrm{vol}(S),\mathrm{vol}(V\setminus S))}.$
-A lower conductance value indicates a more clearly defined community, signifying few edges connecting the set to the rest of the graph relative to its internal connectivity.
-3 The EDQC Algorithm
-This section details the proposed Energy Diffusion-based Quasi-Clique discovery (EDQC) algorithm.
-Given a graph $G=(V,E)$ and a density threshold $\gamma$, EDQC identifies a large vertex set $S$ that induces a $\gamma$-quasi-clique.
-The overall pipeline is outlined in Algorithm 1.
-The algorithm iterates through vertices, treating each as a source to initiate a two-stage process.
-First, an adaptive energy diffusion process generates an energy distribution over the graph.
-Second, guided by this energy distribution, a principled extraction mechanism identifies a high-quality quasi-clique.
-EDQC maintains the largest quasi-clique found across all source vertices.
-Processing vertices in non-increasing degree order is an efficiency heuristic rossi2014fast , as it ensures high-degree vertices are processed as energy sources first, thereby increasing the likelihood of finding a large quasi-clique early in the search.
+A smaller $\Phi(S)$ indicates better separation.
+3. The EDQC Framework
+This section presents EDQC for MQCP under an explicit density threshold $\gamma$.
+We use EDQC to denote the proposed framework.
+Algorithm 1 provides an instantiation of this framework (referred to as the EDQC algorithm when discussing pseudocode and complexity).
+EDQC iterates over vertices as diffusion sources in non-increasing order of degree and maintains the best feasible solution found across all processed sources (lines 2–6).
+This ordering follows a standard efficiency heuristic (Rossi et al., 2014) that prioritizes high-degree sources, which tend to yield larger candidate regions early and thus increase the chance of finding a large feasible quasi-clique in fewer iterations.
+For each source vertex $v$, EDQC first computes an energy score $f:V\to\mathbb{R}_{\geq 0}$ via an adaptive diffusion procedure, providing a locality-aware ranking of vertices around $v$ (line 3).
+It then converts this ranking into a feasible quasi-clique by extracting a candidate set and performing density-constrained refinement to ensure the output satisfies the target threshold $\gamma$ (line 4).
+After computing $S$ for the current source, EDQC keeps it if it improves the best solution so far and returns the best $S^{*}$ at the end (lines 5–7).
 Input: Graph $G=(V,E)$, density threshold $\gamma$, diffusion rounds $T$, activation threshold $\theta$
@@
 […]
-3.1 Adaptive Energy Diffusion
-The first stage, detailed in Algorithm 2, simulates the diffusion of energy from a source vertex $v$ over $T$ synchronous rounds.
-The goal is to produce an energy function $f:V\to\mathbb{R}_{\geq 0}$ where energy is concentrated in structurally cohesive regions surrounding the source.
-The diffusion process is designed to be adaptive.
-The fraction of energy a vertex releases depends on its current energy relative to its neighborhood, a dynamic governed by an adaptive fraction $\alpha(u)$.
-This design is deliberate and distinct from standard diffusion mechanisms such as the heat kernel chung2007heat or PageRank page1999pagerank , which often treat connections uniformly.
-Our adaptive fraction makes the diffusion process sensitive to the evolving energy landscape, enabling EDQC to be more conservative in sparse regions and more aggressive within dense ones.
-The diffusion starts by initializing a unit energy at the source vertex $v$ (line 1).
-In each of the $T$ synchronous rounds (lines 2–14), the algorithm first identifies the set of active vertices, $A$, based on the activation threshold $\theta$ (line 3).
-To ensure updates are synchronous, a temporary copy of the energy function, $f^{\prime}$, is created (line 4).
-The diffusion rule is then applied to each active vertex $u\in A$ with a non-zero degree ($d(u)>0$).
-Isolated vertices are skipped as they have no neighbors to which energy can be transferred, a condition necessary for the conservation of total energy.
-For each such vertex $u\in A$ (line 5), the algorithm computes its adaptive fraction $\alpha(u)$ and the total energy to transfer, $\delta(u)$ (lines 7–8).
-This energy is then stochastically allocated among its neighbors using weights $\phi_{uw}$ derived from an exponential distribution (lines 9–10).
-Subsequently, the energies of the active vertex and its neighbors are updated in the temporary function $f^{\prime}$ (lines 11–13).
-After all active vertices have been processed, the energy function $f$ is updated with the temporary copy, completing the round (line 14).
+3.1. Adaptive Energy Diffusion
+The procedure (Algorithm 2) diffuses a unit amount of energy from a source vertex $v$ for $T$ synchronous rounds.
+It outputs an energy function $f:V\to\mathbb{R}_{\geq 0}$ that induces a locality-aware ranking around $v$.
+The ranking is used for candidate localization, while the density constraint $\gamma$ is enforced later by extraction and refinement.
 Input: Graph $G=(V,E)$, source vertex $v$, diffusion rounds $T$, activation threshold $\theta$
@@
 […]
+Design rationale.
+EDQC runs a fixed-step local diffusion for $T$ rounds and uses two simple mechanisms.
+A vertex is active if its current energy exceeds $\theta$, and the active set is $A=\{x\in V\mid f(x)>\theta\}$ (line 3).
+First, the released amount from an active vertex is adaptive: for $u\in A$, the algorithm releases $\delta(u)=\alpha(u)f(u)$, where $\alpha(u)$ increases with the total energy in the neighborhood of $u$ (line 7–8).
+This adaptivity encourages faster energy exchange when the diffusion remains inside a well-connected region around the source and slows down propagation when the mass reaches low-energy surroundings, which is consistent with diffusion-based local clustering using short-range propagation (Andersen et al., 2006; Kloster and Gleich, 2014; Chung, 2007).
+Second, in each round, EDQC updates only active vertices (line 5), which keeps the computation local and avoids spending time on vertices with negligible energy.
+Compared to classical diffusion models such as PageRank (Page et al., 1999) and the heat kernel (Chung, 2007), EDQC does not aim to approximate a stationary distribution.
+Instead, it runs for a fixed $T$ rounds and uses the scores to induce a local ranking around the source.
+Procedure.
+Algorithm 2 initializes $f(v)=1$ and $f(x)=0$ for $x\neq v$ (line 1).
+In each round, it forms the active set $A$ (line 3) and uses a temporary copy $f^{\prime}$ to ensure synchronicity across updates (line 4).
+For each active vertex $u$ with $d(u)>0$, it computes the adaptive fraction $\alpha(u)$ and the released amount $\delta(u)=\alpha(u)f(u)$ (lines 5–8).
+It then stochastically allocates $\delta(u)$ to neighbors by sampling i.i.d.
+exponential variables and normalizing them into weights $\phi_{uw}$ (line 9–10), and updates $f^{\prime}$ for $u$ and its neighbors accordingly (line 11–13).
+Vertices with $d(u)=0$ have no neighbors to receive energy and are therefore skipped, leaving their energy unchanged.
+Each round ends by setting $f\leftarrow f^{\prime}$ (line 14), and after $T$ rounds the algorithm returns the final energy function $f$ (line 15).
 Properties.
-This diffusion process has several key properties.
+We summarize several basic properties of the diffusion procedure.
 Proposition 0 (Energy Conservation).
@@
 Consider a fixed round and an arbitrary active vertex $u$ where $d(u)>0$.
-The synchronous updates associated with $u$ results in a change of $\delta(u)$ for $u$ and an addition of $\delta(u)\phi_{uw}$ for each neighbor $w\in N(u)$, satisfying $\sum_{w\in N(u)}\phi_{uw}=1$.
+The synchronous updates associated with $u$ result in a change of $\delta(u)$ for $u$ and an addition of $\delta(u)\phi_{uw}$ for each neighbor $w\in N(u)$, satisfying $\sum_{w\in N(u)}\phi_{uw}=1$.
 Therefore, $\delta(u)=\sum_{w\in N(u)}\delta(u)\phi_{uw}$.
 Summing over all active vertices leads to the conclusion that $\sum_{x\in V}f^{\prime}(x)=\sum_{x\in V}f(x)$.
-Here, we should note that the energy of vertices for which $d(u)=0$ or $f(u)=0$ remains unchanged in the energy diffusion process.
+Vertices that are not updated in this round (including those with $f(u)\leq\theta$ or $d(u)=0$) remain unchanged and thus do not affect the sum.
 ∎
-The conservation of energy is a crucial property for the algorithm’s stability.
-It guarantees that the process only redistributes the initial energy, rather than arbitrarily creating or losing it, thus ensuring a fair basis for comparing the energy distributions originating from different source vertices.
+The conservation of energy is a crucial property for the algorithm’s stability. It guarantees that the process only redistributes the initial energy, rather than arbitrarily creating or losing it.
 Proposition 0 (Properties of $\alpha(u)$).
@@
 This confirms that $\alpha(u)$ is strictly increasing with $S(u)$. ∎
-The adaptive nature of our diffusion is formally established by this proposition.
-The bounds on $\alpha(u)$ guarantee a process that is both consistently active and stable, preventing the diffusion from stalling.
-Concurrently, the monotonicity is the key mechanism driving the algorithm’s core premise: it mathematically ensures that energy is exchanged more readily within already high-energy (and thus likely dense) regions, accelerating the concentration of energy.
+The adaptive nature of the diffusion is formalized by this proposition.
+The bounds on $\alpha(u)$ imply that each active vertex always releases at least half, but strictly less than all, of its current energy.
+Thus, the process continues to make progress in every round without becoming unstable.
+Moreover, the monotonic dependence of $\alpha(u)$ on the neighbors’ total energy means that vertices in high-energy neighborhoods release a larger fraction of energy, so diffusion proceeds more rapidly within dense regions and more slowly outside these dense regions.
 Proposition 0 (Unbiased Distribution in Expectation).
-The expected fraction of energy that an active vertex $u$ transfers to a specific neighbor $w\in N(u)$ is uniform across all neighbors: $\mathbb{E}[\phi_{uw}]=1/d(u)$.
+The expected fraction of energy that an active vertex $u$ transfers to a specific neighbor $w\in N(u)$ is uniform across all neighbors, that is, $\mathbb{E}[\phi_{uw}]=1/d(u)$.
 Proof.
@@
 ∎
-While the exponential sampling introduces stochasticity to explore different structural paths, this proposition guarantees that the process remains unbiased in expectation.
-On average, all neighbors are treated fairly, ensuring that the algorithm itself introduces no arbitrary structural bias during the diffusion.
+Conditioned on the current energy $f$, for an active vertex $u$ and any neighbor $w\in N(u)$, the expected update satisfies $\mathbb{E}[f^{\prime}(w)]=f(w)+\delta(u)/d(u)$.
+Thus, the diffusion is uniform over neighbors in expectation; any energy concentration arises from repeated exchanges along the graph structure (as in short-range diffusion for local clustering (Andersen et al., 2006; Kloster and Gleich, 2014; Chung, 2007)).
 Theorem 4 (Complexity of Algorithm 2).
-The EnergyDiffusion algorithm runs in $O(T(n+m))$ time.
+EnergyDiffusion runs in $O(T(n+m))$ time and uses $O(n+m)$ space.
 Proof.
-The algorithm performs $T$ rounds.
-We analyze the computational cost of a single round.
-First, identifying the active set $A$ requires a full scan of all vertices, which takes $O(n)$ time.
-Next, the algorithm iterates through each active vertex $u\in A$.
-For each such vertex, all calculations—including the summations for $\alpha(u)$ and $\phi_{uw}$, and the subsequent updates—are proportional to its degree, $d(u)$.
-Therefore, the total work for processing all active vertices in a round is bounded by the sum of degrees of all vertices in the graph, which is $\sum_{u\in V}O(d(u))=O(m)$.
-Summing the costs within a single round gives a complexity of $O(n+m)$.
-As the process repeats for $T$ rounds, the total complexity of the algorithm is $O(T(n+m))$.
+The algorithm executes $T$ rounds.
+In each round, identifying the active set $A$ incurs a linear scan cost of $O(n)$.
+The subsequent diffusion updates iterate over active vertices, and for each $u\in A$ the work is proportional to $d(u)$.
+Thus the total update cost in one round is $O(\sum_{u\in A}d(u))\leq O(\sum_{u\in V}d(u))=O(m)$.
+Consequently, each round takes $O(n+m)$ time, yielding a total time complexity of $O(T(n+m))$.
+The space usage is $O(n+m)$, dominated by the graph storage and the two auxiliary energy functions $f$ and $f^{\prime}$.
 ∎
-3.2 Quasi-Clique Extraction
-The second stage of EDQC takes the final energy function $f$ returned by Algorithm 2 and produces a vertex set $S$ that induces a $\gamma$-quasi-clique.
-This extraction process (Algorithm 3) has three main operations: an initial ranking of vertices, the identification of a high-quality initial subgraph, and a refinement procedure.
-First, the algorithm identifies the set of active vertices $A$ and sorts them in descending order of their energy scores to create an ordered sequence $P$ (lines 1–2).
-This ranking is crucial, as it places vertices most likely to belong to a cohesive community at the forefront.
-Next, to identify a high-quality initial candidate set, the algorithm performs a sweep cut andersen2006local on the energy-ranked list $P$ (line 3).
-This procedure examines the nested prefixes of the list and selects the one that minimizes conductance, as it represents the most well-defined community structure.
-This strategy is highly effective because the initial energy ranking provides a high-quality vertex permutation.
-It is a known principle in local graph clustering that sweep cuts perform best when applied to an ordering that places vertices from the same community contiguously andersen2006local .
-Our energy diffusion process is designed for precisely this purpose: it creates an ordering where structurally cohesive vertices are naturally placed at the top.
-This, in turn, enables the sweep cut to accurately identify the boundary of the best candidate subgraph.
-Finally, this initial set $S$ undergoes a refinement procedure (lines 4–15).
-The refinement begins with a pruning loop (lines 4–6).
-If the initial set $S$ does not meet the density threshold, this loop iteratively removes the vertex with the minimum internal degree $d_{S}(u)$ until the set $S$ induces a valid $\gamma$-quasi-clique.
-Following this, an expansion phase seeks to enlarge the set (lines 8–15).
-In each pass, the remaining candidates $R$ are re-sorted based on their connectivity to the current set $S$ (line 10).
-The algorithm then iterates through this sorted list $R$, attempting to greedily add the first vertex $v$ that maintains the density constraint (lines 12–13).
-Upon a successful insertion, the current iteration over the candidate list is immediately terminated (line 15), allowing the algorithm to restart the process with an up-to-date candidate ranking in the next iteration.
+3.2. Quasi-Clique Extraction
+The procedure (Algorithm 3) converts the energy function $f$ into a $\gamma$-quasi-clique by first ranking active vertices by energy, then selecting an initial candidate via a conductance sweep cut over energy-ranked prefixes, and finally refining the candidate to satisfy the density threshold $\gamma$.
+Specifically, the sweep cut selects the minimum-conductance prefix to turn the ranking into an explicit candidate region.
+The refinement step then enforces feasibility under $\gamma$ through pruning and greedy expansion.
 Input: Graph $G=(V,E)$, energy function $f$, density threshold $\gamma$, activation threshold $\theta$
@@
 3 $P\leftarrow$ vertices of AA sorted by decreasing $f$;
-4
+4 if $|P|<2$ then return $\emptyset$;
+5
 […]
-6
-7while $|S|>1$ and $\eta(G[S])<\gamma$ do
+7 $S\leftarrow\{v_{1},\dots,v_{i^{*}}\}$;
+8
+9while $|S|>1$ and $\eta(G[S])<\gamma$ do
 10  $u\leftarrow\arg\min_{x\in S}d_{S}(x)$;
 11  $S\leftarrow S\setminus\{u\}$;
-10
-11
+12
+13
+14if $|S|<2$ then return $\emptyset$;
+15
 […]
@@
 24        break;
-21
-22
-23
+25
+26
+27
 return $S$
 […]
+First, the algorithm collects active vertices $A$ and sorts them by decreasing energy to form an ordered list $P$ (lines 1–2).
+If $|P|<2$, it returns $\emptyset$ (line 3).
+This ranking is important, as it places vertices that are more likely to belong to a cohesive region near the front of the sequence.
+Next, to obtain an initial candidate set, the algorithm performs a sweep cut (Andersen et al., 2006) on the energy-ranked list $P$ (lines 4–5).
+Formally, let $P=(v_{1},\dots,v_{k})$.
+The procedure evaluates the conductance of each prefix $S_{i}=\{v_{1},\dots,v_{i}\}$ for $1\leq i\leq k-1$, and selects the minimum-conductance prefix as the initial set $S$.
+Note that we evaluate prefixes up to k1k-1 to avoid the degenerate case $S=V$ (or $S=P$ here), for which $\Phi(S)$ is undefined as the denominator becomes zero.
+In this way, the diffusion ranking is converted into a vertex set whose boundary is identified by the conductance metric.
+Since $P$ is induced by diffusion scores, low-conductance prefixes are natural candidates for subsequent refinement.
+Finally, the resulting initial set $S$ undergoes a refinement procedure (lines 6–18).
+The refinement begins with a pruning loop (lines 6–8).
+If the pruning reduces $S$ to a size smaller than $2$, the procedure returns an empty set (line 9), since a $\gamma$-quasi-clique requires at least two vertices.
+If the initial set $S$ does not meet the density threshold, this loop iteratively removes the vertex with the minimum internal degree $d_{S}(u)$ until the set $S$ induces a valid $\gamma$-quasi-clique.
+Following this, an expansion phase seeks to enlarge the set while preserving feasibility (lines 10–18).
+In each pass, the remaining candidates $R=P\setminus S$ are re-sorted based on their connections to the current set $S$ (line 13).
+The algorithm then iterates through $R$ and greedily adds the first vertex $v$ that maintains the density constraint (lines 15–16).
+Upon a successful insertion, the current iteration over $R$ is terminated, since the connectivity ordering depends on the updated set $S$.
+The process repeats until no further insertion is possible without violating $\gamma$.
 Theorem 5 (Complexity of Algorithm 3).
-Let $k=|A|$ be the number of active vertices. The ExtractQuasiClique stage runs in $O(k^{2}+m(A))$, where $m(A)$ is the number of edges in the induced subgraph $G[A]$.
+Let $k=|A|$. ExtractQuasiClique runs in $O(k^{2}+\mathrm{vol}(A))$ time and uses $O(k)$ space.
 Proof.
-The overall complexity is determined by the main steps of the algorithm.
-Sorting the $k$ active vertices takes $O(k\log k)$ time.
-The subsequent sweep cut, which involves calculating conductance for $k$ prefixes, can be implemented in $O(m(A))$ time with incremental updates for the cut and volume.
-The final refinement stage is the most computationally intensive.
-The pruning and expansion loops can, in the worst case, lead to a series of $O(k)$ vertex additions or removals.
-With incremental maintenance of internal degrees and edge counts, each step of these loops can take up to $O(k)$ time to find the best candidate to remove or to scan through remaining candidates for addition.
-This results in a worst-case complexity of $O(k^{2})$ for the refinement phase.
-Summing these costs, the total complexity is $O(k\log k+m(A)+k^{2})$, which simplifies to $O(k^{2}+m(A))$.
+Sorting the $k$ active vertices takes $O(k\log k)$.
+The sweep cut linearly scans the sorted list; computing the conductance of all prefixes via incremental updates takes $O(\mathrm{vol}(A))$ time.
+The refinement stage entails a greedy optimization sequence.
+In the worst case, it performs $O(k)$ vertex additions or removals.
+By maintaining internal degrees incrementally, each update requires $O(k)$ operations, yielding an $O(k^{2})$ bound for refinement.
+Since $k\log k\in O(k^{2})$, the total time complexity is $O(k^{2}+\mathrm{vol}(A))$.
+The space usage is dominated by the ranked list and auxiliary structures for vertex membership and internal degrees, which use $O(k)$ space.
 ∎
 Theorem 6 (Complexity of EDQC).
-The theoretical worst-case time complexity of the EDQC algorithm (Algorithm 1) is $O(n^{3})$.
+Fix a source vertex $v$.
+Let $k=|A|$.
+One iteration of EDQC with source $v$ runs in $O\!\big(T(k+\mathrm{vol}(A))+k^{2}\big)$ time and uses $O(n+m)$ space.
+The total time complexity for processing $B$ sources is the summation of this bound over the $B$ iterations.
+In the degenerate worst case where $B=n$ and $k=\Theta(n)$ in every iteration, this bound recovers $O(n^{3})$.
 Proof.
-The algorithm’s main loop iterates up to $n$ times.
-Each iteration calls two procedures: EnergyDiffusion, with a complexity of $O(T(n+m))$, and ExtractQuasiClique, with a worst-case complexity of $O(n^{2}+m)$.
-As $T$ is a small constant, the cost of a single iteration is dominated by the extraction step, at $O(n^{2}+m)$.
-Therefore, the total worst-case complexity is $O(n^{3}+nm)$.
-Since $m=O(n^{2})$ for any simple graph, this simplifies to $O(n^{3})$.
-The initial sorting of vertices by non-increasing degree is a lower-order term that does not affect this bound.
-We note that this is a theoretical worst-case analysis. In practice, EDQC is run as a time-bounded heuristic, and its empirical efficiency is demonstrated in our experimental section. ∎
-4 Experimental Evaluation
-We conducted a comprehensive experimental evaluation to validate the performance and robustness of EDQC.
-This section is organized as follows.
-We first detail the experimental setup, the baseline methods, and the datasets used.
-We then present the main results on effectiveness and efficiency, followed by a statistical analysis to confirm the significance of our findings.
-Subsequently, we provide an in-depth analysis of the algorithm, including an ablation study to verify the contribution of its key components and an empirical study to validate its core premise.
-The section concludes with a summary of our parameter sensitivity analysis.
-4.1 Experimental Setup
-All methods were implemented in C++ and compiled with g++ 9.4.0 using the ‘-O3’ optimization flag.
-The experiments were conducted on a machine equipped with an AMD EPYC 7H12 CPU (1.5 GHz base frequency) and 128 GB of RAM, running Ubuntu 20.04.
-Each baseline was run with the parameter settings recommended in its original paper.
-For EDQC, we fixed the diffusion rounds at $T=2$ and the activation threshold at $\theta=10^{-4}$.
-These values were selected based on a sensitivity study, the details of which are provided in Section 4.8.
+Consider a fixed source $v$.
+The diffusion stage runs for $T$ rounds.
+In our implementation, the active set is maintained incrementally, so each round updates only active vertices and scans their incident edges, incurring $O\!\left(T(k+\mathrm{vol}(A))\right)$ time.
+The extraction stage sorts the $k$ active vertices in $O(k\log k)$.
+The subsequent sweep cut requires $O(\mathrm{vol}(A))$ time via incremental updates; since $\mathrm{vol}(A)\leq T(k+\mathrm{vol}(A))$ for $T\geq 1$, this cost is subsumed by the diffusion stage.
+The refinement stage performs a greedy optimization sequence; with incremental degree maintenance, its worst-case cost is bounded by $O(k^{2})$.
+Therefore, the total time complexity of one iteration is $O\!\big(T(k+\mathrm{vol}(A))+k^{2}\big)$; in the degenerate worst case where $k=\Theta(n)$ and $B=n$, the total time over all sources becomes $O(n^{3})$.
+The space usage is $O(n+m)$, dominated by storing $G$ and two reusable energy functions $f$ and $f^{\prime}$.
+∎
+4. Experimental Evaluation
+We conducted a comprehensive experimental evaluation of EDQC.
+We report solution quality, stability, and runtime under instance-specific density thresholds, together with statistical tests, ablations, and additional analyses on robustness across $\gamma$ and the energy–density premise.
+4.1. Experimental Setup
+All algorithms were implemented in C++ and compiled with g++ 9.4.0 using the “-O3” optimization flag.
+Experiments were run on a machine with an AMD EPYC 7H12 CPU (1.5 GHz base frequency), 128 GB of RAM, and Ubuntu 20.04.
+Baseline parameters followed the original papers.
+For EDQC, we fixed $T=2$ and $\theta=10^{-4}$ (chosen based on the sensitivity study in Section 4.10) across all datasets.
 Due to inherent randomness, all non-deterministic methods (all except NBSim) were executed 10 times with random seeds from 1 to 10; NBSim, being deterministic, was run once.
-For the non-deterministic methods, we report their average quasi-clique size and standard deviation (STD) over the 10 runs, presented in the format ‘average size $\pm$ STD’, where the standard deviation is reported only if it is non-zero.
+For the non-deterministic methods, we report their average quasi-clique size and standard deviation (STD) over the 10 runs, presented in the format “average size $\pm$ STD”, where the standard deviation is reported only if it is non-zero.
 All runs were subject to a 60-second timeout.
-If an algorithm failed to find any valid $\gamma$-quasi-clique in all 10 runs for an instance, we mark its result as ‘N/A’.
-The 60-second timeout reflects common practice in recent time-bounded heuristic studies lin2017reduction ; zheng2023farsighted .
-4.2 Baseline Methods
+If an algorithm failed to find any valid $\gamma$-quasi-clique in all 10 runs for an instance, we mark its result as “N/A”.
+The 60-second timeout reflects common practice in recent time-bounded heuristic studies (Lin et al., 2017; Zheng et al., 2023).
+We report wall-clock time excluding file I/O, and run all methods in a single thread.
+4.2. Baseline Methods
 We compared EDQC against four state-of-the-art algorithms:
-  • NBSim/FastNBSim pang2024similarity : Similarity-based heuristics that expand subgraphs based on neighborhood overlap. FastNBSim accelerates NBSim via MinHash-based approximation.
-  • NuQClqF1/NuQClqR1 liu2024optimization : Advanced variants of NuQClq chen2021nuqclq .
+  • NBSim/FastNBSim (Pang et al., 2024): Similarity-based heuristics that expand subgraphs based on neighborhood overlap. FastNBSim accelerates NBSim via MinHash-based approximation.
+  • NuQClqF1/NuQClqR1 (Liu et al., 2024): Advanced variants of NuQClq (Chen et al., 2021).
 These algorithms employ information feedback models that integrate historical solutions using fitness-weighted scores to guide local search.
-4.3 Datasets
-We evaluated all algorithms on 75 real-world datasets from SNAP snapnets and the Network Repository nr .
+4.3. Datasets
+We evaluated all algorithms on 75 real-world datasets from SNAP (Leskovec and Krevl, 2014) and the Network Repository (Rossi and Ahmed, 2015).
 The collection covers diverse domains, including web, social, and communication graphs.
 The datasets span a wide spectrum of scales, with vertex counts ($n$) ranging from 62 to over 16.7 million, and edge densities spanning from $8.41\times 10^{-2}$ down to as low as $1.87\times 10^{-7}$.
-This diversity provided a challenging testbed for evaluating the scalability and effectiveness of the algorithms.
-4.4 Main Results
-We now present the primary results comparing the effectiveness and efficiency of EDQC against the baselines.
-A key challenge in this comparison is that NBSim and FastNBSim do not accept a density threshold $\gamma$ as input.
-To facilitate a fair evaluation, we designed two separate experimental settings (Table 1 and Table 2).
-To save space, we omit instances where all four compared algorithms found solutions of the same average size.
-Table 1: Quasi-clique sizes with the threshold $\gamma$ set by NBSim’s output.
+4.4. Main Results
+This section reports the main results on solution quality, stability, and runtime under instance-specific density thresholds induced by NBSim and FastNBSim.
+A key challenge in this comparison is that NBSim and FastNBSim do not take a density threshold $\gamma$ as input, whereas EDQC, NuQClqF1, and NuQClqR1 do.
+To enable a comparable evaluation across these methods, we adopt two evaluation settings, following the protocol of Pang et al.
+(Pang et al., 2024).
+Table 1. Quasi-clique sizes with the threshold $\gamma$ set by NBSim’s output.
 Instance $\gamma$ NBSim NuQClqF1 NuQClqR1 EDQC
@@
 rgg_n_2_24_s0 0.98 20 19.7 $\pm$ 0.95 18.7 $\pm$ 0.48 23
-Table 2: Quasi-clique sizes with $\gamma$ set by FastNBSim’s max density over 10 runs. For FastNBSim, stats are on successful runs only (rate shown in parentheses).
+In the first setting, NBSim is run once per instance, and the density of its output quasi-clique is used as the target threshold $\gamma$ for EDQC, NuQClqF1, and NuQClqR1.
+In the second setting, FastNBSim is run 10 times following the protocol in Section 4.1; the density threshold $\gamma$ is then set to the maximum density achieved over these runs, which reflects the strongest cohesion level FastNBSim can reach on the instance, and EDQC, NuQClqF1, and NuQClqR1 are evaluated under this common threshold.
+The detailed quasi-clique sizes in these two settings are reported in Tables 1 and 2; to save space, instances where all algorithms obtain quasi-cliques of identical average size are omitted.
+Table 2. Quasi-clique sizes with $\gamma$ set by FastNBSim’s max density over 10 runs. For FastNBSim, stats are on successful runs only (rate shown in parentheses).
 Instance $\gamma$ FastNBSim NuQClqF1 NuQClqR1 EDQC
@@
 Beyond average size, EDQC demonstrates exceptional stability.
-It consistently reports a standard deviation close to zero (e.g., $0.32$ on the large eu-2005 graph), even on complex instances.
+It consistently reports a standard deviation close to zero (e.g., $0.32$ on the large eu-2005 graph under the induced $\gamma$), even on complex instances.
 This is in stark contrast to the NuQClq variants, which frequently exhibit high variance on the same instances (e.g., a STD of $115.84$ for NuQClqF1 on eu-2005).
@@
 This lack of explicit density control makes them less suitable for applications requiring highly cohesive structures, a task where EDQC’s reliability is a distinct advantage.
-Figure 2: Runtime comparison of EDQC against baselines on a log-log scale.
 Efficiency.
-The efficiency of EDQC is compared against baselines in Figure 2, where, following prior work sunnumds , instances where the runtime of both algorithms is below 0.1 seconds are excluded.
+The efficiency of EDQC is compared against baselines in Figure 2, where, following prior work (Sun et al., 2025), instances where the runtime of both EDQC and the corresponding baseline is below 0.1 seconds are excluded.
 On this log-log plot, points falling below the $1\times$, $10\times$, and $100\times$ lines indicate that EDQC is faster than the baseline by the corresponding factor or more.
-The results clearly show that EDQC is substantially more efficient than the high-performing local search methods, NuQClqF1 and NuQClqR1, with a large number of points falling below the $10\times$ line, signifying a speedup of over an order of magnitude.
-In comparison to the fast similarity-based heuristics (NBSim and FastNBSim), EDQC’s runtime remains highly competitive, as most of their corresponding points cluster around the $1\times$ line.
-Overall, this demonstrates that EDQC strikes an excellent balance between achieving state-of-the-art solution quality and maintaining high computational efficiency.
-4.5 Statistical Analysis
-(a) Comparison in the NBSim setting.
-(b) Comparison in the FastNBSim setting.
-Figure 3: Critical Difference diagrams for algorithm ranks on all 75 datasets (significance level 0.05). A horizontal bar connects algorithms with no statistically significant difference.
-To assess statistical significance, we applied the Friedman test friedman1937use on the algorithm rankings across all 75 datasets.
-The null hypothesis that all algorithms perform equally was rejected in both settings ($p<0.05$), allowing us to proceed with the post-hoc Nemenyi test demvsar2006statistical to identify pairwise performance differences.
+The results indicate that EDQC is substantially more efficient than the local search methods, NuQClqF1 and NuQClqR1, with many points falling below the $10\times$ line, signifying a speedup of over an order of magnitude.
+In comparison to the similarity-based heuristics (NBSim and FastNBSim), EDQC’s runtime remains competitive, as most of their corresponding points cluster around the $1\times$ line.
+Overall, this demonstrates that EDQC strikes a competitive balance between achieving state-of-the-art solution quality and maintaining high computational efficiency.
+Figure 2. Runtime comparison of EDQC against baselines on a log-log scale.
+4.5. Statistical Analysis
+To assess statistical significance, we applied the Friedman test (Friedman, 1937) on the algorithm rankings across all 75 datasets.
+The null hypothesis that all algorithms perform equally was rejected in both settings ($p<0.05$), allowing us to proceed with the post-hoc Nemenyi test (Demšar, 2006) to identify pairwise performance differences.
+On each dataset, we rank algorithms by the reported quasi-clique size under the same induced $\gamma$ (for randomized methods, the mean size over 10 runs).
+Ties receive averaged ranks.
+When a method is marked as “N/A” on a dataset, we treat it as the worst performance for ranking on that dataset.
 The results are visualized using Critical Difference (CD) diagrams in Figure 3.
@@
 Its superiority is also confirmed to be statistically significant, as it significantly outperforms NuQClqF1, NuQClqR1, and FastNBSim.
-In summary, the statistical tests provide conclusive evidence for the superiority of EDQC.
-In both experimental settings, EDQC not only achieves the best average rank but also demonstrates a statistically significant performance advantage over all compared state-of-the-art baselines.
-These results provide strong statistical evidence that EDQC represents a reliably superior approach for this task.
-4.6 Analysis of EDQC Components
+(a) Comparison in the NBSim setting.
+(b) Comparison in the FastNBSim setting.
+Figure 3. Critical Difference diagrams for algorithm ranks on 75 datasets (significance level 0.05). A horizontal bar connects algorithms with no statistically significant difference.
+In summary, the statistical tests provide strong evidence that EDQC achieves better average ranks than the compared baselines under both experimental settings at the 0.05 significance level.
+4.6. Analysis of EDQC Components
 To isolate and verify the contribution of each key component, we conducted an ablation study comparing the full EDQC against five variants across 112 unique experimental settings (consolidating settings where both configurations yielded the same $\gamma$ value).
 These variants were designed to test the core aspects of our algorithm: the subgraph identification strategy (EDQC-1), the energy-guided ranking (EDQC-2, EDQC-3), the refinement phase (EDQC-4), and the nature of the diffusion process itself (EDQC-5).
-The results, summarized in Table 3, provide conclusive evidence that each component is essential for EDQC’s superior performance.
+The results, summarized in Table 3, provide strong evidence that each component contributes to the overall performance of EDQC.
 The most dramatic performance degradation occurred with EDQC-5, which replaced our adaptive diffusion with a naive random process, and EDQC-1, which substituted conductance-based identification with a max-density heuristic.
@@
 As variants with a fixed processing order for active vertices produced identical results to the full EDQC, the order-independence guaranteed by the synchronous design was empirically verified.
-Table 3: Comparison of the full EDQC with its five simplified variants across 112 experimental settings.
+Table 3.
+Comparison of the full EDQC with its five simplified variants across 112 experimental settings.
 #bet and #wor represent the number of instances where EDQC achieves better and worse quasi-clique sizes, respectively.
@@
 EDQC-5 Naive Random Diffusion 106 0
-Figure 4: Subgraph density vs.
+4.7. Results under Fixed Density Thresholds
+Figure 4.
+Fixed-$\gamma$ comparison of EDQC, NuQClqF1, and NuQClqR1 on 75 graphs for $\gamma\in\{0.7,0.8,0.9,0.95,0.99\}$.
+Left: normalized solution quality (1.0 = best average quasi-clique size at that $\gamma$).
+Right: normalized instability (0 = lowest, 1 = highest average standard deviation at that $\gamma$).
+Beyond the instance-specific thresholds induced by NBSim and FastNBSim, we also examine how algorithms that accept an explicit density parameter behave across a shared set of fixed thresholds.
+Specifically, we compare EDQC, NuQClqF1, and NuQClqR1 on all 75 datasets under $\gamma\in\{0.7,0.8,0.9,0.95,0.99\}$.
+We exclude NBSim and FastNBSim here since they do not take $\gamma$ as input.
+For each graph and each $\gamma\in\{0.7,0.8,0.9,0.95,0.99\}$, the three algorithms are run under the same protocol as in Section 4.1.
+For each dataset and each (algorithm, $\gamma$), we compute the mean and standard deviation (STD) of the quasi-clique size over 10 runs, and then average these quantities over the 75 datasets.
+If an algorithm fails to find any valid $\gamma$-quasi-clique on a dataset, we treat it as the worst performance for aggregation.
+The left panel of Figure 4 shows a radar plot of the normalized solution quality, where for each $\gamma$ the best average size among the three algorithms is scaled to $1.0$ (computed per $\gamma$ across the three algorithms).
+The right panel reports the normalized instability, where for each $\gamma$ the lowest average STD is mapped to 0 and the highest to $1$ (again per $\gamma$).
+We use a radar plot to visualize the performance profile across multiple $\gamma$ values.
+Figure 5.
+Subgraph density vs.
 total retained energy on eight representative datasets.
 Each blue dot represents one of 1,000 randomly sampled subgraphs of a fixed size; the red marks the quasi-clique of the same size found by EDQC.
-4.7 Empirical Validation of the Energy-Density Premise
+The plots reveal a consistent pattern.
+On the solution-quality radar (left), EDQC stays close to $1.0$ across all density thresholds, indicating that it attains the best (or tied-best) average size at each $\gamma$.
+Moreover, the gap in normalized size tends to widen as $\gamma$ increases from $0.7$ to $0.99$, suggesting that EDQC is particularly strong when stricter cohesion is required.
+On the instability radar (right), EDQC remains near 0 across all densities, reflecting consistently low variability, while NuQClqF1 shows higher normalized instability and NuQClqR1 lies in between.
+These fixed-threshold results provide additional evidence that EDQC delivers both high solution quality and low run-to-run variability over a broad range of density requirements.
+4.8. Structural characterization on web-Google
+We further examine the structural properties of the $\gamma$-quasi-cliques returned by EDQC and NuQClqF1/R1 on the large web graph web-Google at $\gamma=0.9$.
+Table 4 reports the mean and standard deviation over 10 runs.
+EDQC finds substantially larger groups ($60.0\pm 0.0$ nodes vs.
+$43.8\pm 8.4$ and $41.2\pm 6.5$) while maintaining similarly high internal density.
+At the same time, its conductance is much lower and far more stable ($0.101\pm 0.003$ vs.
+$0.570\pm 0.305$ and $0.669\pm 0.235$), indicating that EDQC consistently identifies larger and better isolated cohesive groups.
+The clustering coefficients are very high and comparable ($\approx 0.92$$0.93$), suggesting similar internal triangle structure across methods.
+Table 4. Structural properties of $\gamma$-quasi-cliques on web-Google at $\gamma=0.9$. Mean $\pm$ standard deviation over 10 runs.
+Method $|S|$ Density $\Phi(S)$ Clust. coeff.
+EDQC $60.0$ $0.904\pm 0.002$ $0.101\pm 0.003$ $0.921\pm 0.001$
+NuQClqF1 $43.8\pm 8.4$ $0.911\pm 0.010$ $0.570\pm 0.305$ $0.934\pm 0.015$
+NuQClqR1 $41.2\pm 6.5$ $0.907\pm 0.004$ $0.669\pm 0.235$ $0.929\pm 0.015$
+4.9. Empirical Validation of the Energy-Density Premise
 To empirically validate our algorithm’s core premise that energy concentrates in dense regions, we analyzed the relationship between subgraph density and energy retention on eight representative datasets.
@@
 This strong alignment provides the empirical justification for using the energy ranking as a powerful guide in the discovery of quasi-cliques.
-4.8 Parameter Sensitivity
-This section provides the detailed results of our parameter sensitivity analysis, which was designed to evaluate the impact of the number of diffusion rounds $T$ and the activation threshold $\theta$ on EDQC’s performance.
-For this study, we selected three representative datasets to cover different scales: ego-facebook (a small graph, $|V|\leq 10^{4}$), rgg_n_2_19_s0 (a medium graph, $10^{4}<|V|\leq 10^{6}$), and web-Google (a large graph, $|V|>10^{6}$).
-On each of these graphs, we evaluated EDQC by testing 15 combinations of $T\in\{1,2,3\}$ and $\theta\in\{10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3},10^{-2}\}$.
-This evaluation was repeated under three different density thresholds, $\gamma\in\{0.7,0.8,0.9\}$, to ensure a comprehensive analysis.
-We report both the discovered quasi-clique size and the runtime for each configuration.
-The results for the small, medium, and large graphs are presented in Figure 57.
-Figure 5: Parameter sensitivity analysis for EDQC on the small representative graph (ego-facebook).
-Each of the three rows corresponds to a different density threshold $\gamma$.
-For each $\gamma$, the left plot shows the resulting quasi-clique size versus the activation threshold $\theta$ (on a log scale), and the right plot shows the corresponding runtime.
-Figure 6: Parameter sensitivity analysis on the medium representative graph (rgg_n_2_19_s0). The layout of the plots is identical to that described in the caption of Figure 5.
-Figure 7: Parameter sensitivity analysis on the large representative graph (web-Google). The layout of the plots is identical to that described in the caption of Figure 5.
-Across all three datasets, a consistent pattern emerges.
-In terms of solution quality (size), configurations with $T=2$ and $T=3$ consistently and significantly outperform $T=1$, especially for smaller values of the activation threshold ($\theta\leq 10^{-3}$).
-As $\theta$ increases, making vertex activation stricter, the performance of all settings tends to decline.
-Regarding efficiency, the runtime plots clearly show that the computational cost increases with $T$.
-While $T=3$ occasionally finds a marginally larger quasi-clique than $T=2$, its runtime is substantially higher across all settings.
-For instance, in the large graph setting (Figure 7), the runtime for $T=3$ can be several times that of $T=2$ for little to no improvement in solution size.
-Based on this analysis, we conclude that the configuration $T=2$ provides the best trade-off between solution quality and computational cost.
-Paired with a small activation threshold such as $\theta=10^{-4}$, which consistently lies in the stable, high-performance region shown in the plots, this configuration represents an effective and efficient choice.
-This justifies our use of $T=2$ and $\theta=10^{-4}$ for all main experiments presented in the paper.
-5 Conclusion
-In this paper, we introduced EDQC, a novel algorithm for quasi-clique discovery designed to mitigate key limitations of existing methods, such as their sensitivity to random seeds and lack of explicit edge-density guarantees.
-We demonstrated that the effectiveness of EDQC stems from its two-stage design.
-First, an adaptive energy diffusion process generates an energy ranking that highlights structurally cohesive regions.
-Then, guided by this ranking, a principled extraction mechanism based on conductance minimization identifies and refines a high-quality subgraph.
-To the best of our knowledge, EDQC is the first method to incorporate energy diffusion into this task, offering a robust alternative to traditional heuristics.
-Our extensive experiments on 75 real-world graphs confirmed the effectiveness of this approach.
-The results demonstrated that EDQC finds larger quasi-cliques on most datasets, with consistently lower variance across runs, highlighting its superior stability.
-Furthermore, statistical tests confirmed the significance of these findings, and our analysis showed that EDQC maintains a competitive runtime, striking an excellent balance between solution quality and efficiency.
-Future work can extend the EDQC framework in several promising directions.
-Generalizing the diffusion model for weighted and attributed graphs is an important next step.
-Applying EDQC to downstream tasks, such as anomaly detection in financial networks and community detection in social graphs, could further validate its practical utility.
-Finally, exploring more advanced refinement strategies may lead to further improvements in solution quality.
+4.10. Parameter Sensitivity
+We studied the sensitivity of EDQC to the diffusion rounds $T$ and the activation threshold $\theta$ on three datasets of different scales (ego-facebook, rgg_n_2_19_s0, and web-Google).
+We tested $T\in\{1,2,3\}$ and $\theta\in\{10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3},10^{-2}\}$ under three values of $\gamma$.
+Due to space limits, Figure 6 reports results on web-Google.
+Overall, EDQC is stable over a broad range of $(T,\theta)$.
+Increasing $T$ from $1$ to $2$ typically improves solution quality, while $T=3$ offers little additional gain but noticeably increases runtime.
+Smaller thresholds ($10^{-4}$ to $10^{-3}$) yield similar solution quality, whereas larger values ($\theta\geq 5\times 10^{-3}$) tend to over-prune and slightly degrade performance, especially at higher $\gamma$.
+Figure 6. Sensitivity of EDQC to $(T,\theta)$ on web-Google under three density thresholds $\gamma$. Left: quasi-clique size; right: runtime (log-scale $\theta$).
+5. Conclusion
+This paper proposes EDQC, an effective framework for cohesive group discovery under explicit edge-density constraints.
+EDQC leverages a lightweight adaptive energy diffusion process to rank vertices and localize candidate regions in interaction graphs; guided by this ranking, it performs conductance-based extraction and density-feasible refinement to obtain a $\gamma$-quasi-clique that satisfies the target threshold.
+This design aims to alleviate practical limitations observed in existing heuristics, including sensitivity to randomness and the lack of explicit density guarantees.
+Extensive experiments on 75 real-world graphs across varying density thresholds show that EDQC identifies larger mean $\gamma$-quasi-cliques in the vast majority of cases, with consistently lower variance across runs while maintaining competitive runtime, and statistical tests support the significance of these improvements.
+Overall, EDQC offers a robust primitive for cohesive group discovery in interaction graphs, aligning with graph-based recommender systems where explicitly constrained cohesive groups underpin tasks such as social recommendation, bundle discovery, and community-aware modeling.
+Future work includes extending EDQC to weighted and attributed graphs and integrating it into end-to-end recommender pipelines to quantify task-level impact.